### v1

# Degree-one foliations on complete intersections

Abstract.
We prove that, under mild restrictions, the space of codimension-one foliations of degree one on a smooth projective complete intersection has two irreducible components of logarithmic type. We also prove that the same conclusion holds for any smooth hypersurface of dimension at least three that is not a quadric threefold. The proof of these results follows essentially from a more general structure theorem for foliations on manifolds covered by lines.
Key words and phrases:
holomorphic foliations, holomorphic distributions, complete intersections, hypersurfaces, extensions, restrictions
2020 Mathematics Subject Classification:
Primary: 37F75, 32S65; Secondary: 14M25
Figueira was partially supported by a postdoctoral fellowship from the CAPES-PIBID program in Mathematics at the Universidade Federal Fluminense (UFF). Kuster was supported by CNPq Grant Number 140986/2021-9 and CAPES - PDSE - Programa de Doutorado Sanduíche no Exterior Grant Number 88881.846472/2023-01. Lizarbe was supported by CNPq/MCTI/FNDCT No. 18/2021 (“Global geometry of connections, holomorphic webs and foliations,” Grant Number 402936/2021-3) and also by CNPq/MCTI No. 10/2023 – Edital Universal (“Geometria das equações diferenciais algébricas,” Grant Number 408687/2023-1).
1. Introduction
A singular holomorphic foliation $\mathcal{F}$ of codimension-one on a projective variety $X$ is determined by a $1$-form $\omega$ with values in a line bundle $N$, satisfying the Frobenius integrability condition $\omega\wedge d\omega=0$, and whose singular set has codimension at least two. In this case, $N$ is called the normal bundle of $\mathcal{F}$. The set of all such foliations with fixed normal bundle $N$ defines a quasi-projective variety, denoted by ${\sf Fol}(X,N)$, known as the space of codimension-one foliations on $X$ with normal bundle $N$.
The irreducible components of these spaces have been studied in various contexts. When $\operatorname{Pic}(X)\simeq\mathbb{Z}$, a foliation $\mathcal{F}$ with normal bundle $N$ is said to have degree $d$ if $N\simeq{\mathcal{O}}_{X}(d+2)$, where ${\mathcal{O}}_{X}(1)$ denotes the hyperplane bundle. In this case, we denote the space of degree-$d$ foliations by ${\sf Fol}(X,d)$. For example, when $X$ is a surface, the integrability condition is automatically satisfied, and the corresponding spaces are irreducible. When $X={\mathbb{P}^{n}}$, for $n\geq 3$, the cases $d=0,1,2$ are completely classified, and partial results exist for $d=3$ (see [13], [16], [7], and [10]). For more general smooth projective varieties, few results describe the structure of foliations of low degree. When $d=0$, classifications are available for certain homogeneous varieties [5, 19], for hypersurfaces [15], and for smooth complete intersections [1].
In this work, we study the irreducible components of the space of degree-one foliations on smooth projective complete intersections. Our main result is the following:
Theorem A.
Let $X$ be a smooth complete intersection in ${\mathbb{P}^{n+s}}$ of type $(d_{1},\ldots,d_{s})$, with $n\geq 4$ and $c_{1}(TX)\geq 2$. If either $s=1$ (i.e., $X$ is a hypersurface) or $d_{j}\geq 3$ for all $j=1,\dots,s$, then ${\sf Fol}(X,1)$ has exactly two irreducible components.
The general elements in these components are defined by logarithmic forms (see §2.2). In particular, all foliations in these spaces arise as restrictions of foliations on the ambient projective space (see Proposition 3.1). When $X$ is the smooth three-dimensional quadric, it is shown in [20] that ${\sf Fol}(X,1)$ possesses an additional irreducible component, consisting of rigid foliations arising from actions of the affine Lie algebra on the quadric, totaling three components.
Among hypersurfaces, the quadric is the only exceptional case. Theorem A applies to all smooth hypersurfaces of degree $\deg(X)\leq n$ and dimension $n\geq 4$. Moreover, by [15, Theorem 3.6], every degree-one foliation on a hypersurface $X$ extends to ${\mathbb{P}^{n+1}}$ when $n\geq 3$ and $\deg(X)\geq 4$. We complement this result by showing that it also holds for cubic threefolds:
Theorem B.
Let $X$ be a smooth hypersurface in ${\mathbb{P}^{n+1}}$ with $n\geq 3$ and $(n,\deg(X))\neq(3,2)$. Then ${\sf Fol}(X,1)$ has exactly two irreducible components. Every such foliation is the restriction of a foliation on ${\mathbb{P}^{n+1}}$.
To prove Theorem A, we focus on varieties covered by lines. We study the deformations of lines along the leaves of a foliation, as introduced in [21]. First, we show that a general free line is not invariant under any foliation (see Proposition 4.7).
A detailed analysis of the tangential foliation leads to the following structure theorem for foliations of small degrees on varieties covered by lines (see Theorem 4.8):
Theorem C.
Let $X\subset{\mathbb{P}^{N}}$ be a smooth projective variety of dimension $n\geq 3$ with $\operatorname{Pic}(X)=\mathbb{Z}$, covered by lines. If $\mathcal{F}$ is a codimension-one singular holomorphic foliation on $X$ of degree $d$ such that either $d=1$ or
$c_{1}(TX)\geq n+d-2,$
then one of the following holds:
  1. (1)
    $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros;
  2. (2)
    $\mathcal{F}$ is algebraically integrable;
  3. (3)
    $\mathcal{F}$ admits an algebraically integrable subfoliation of codimension two and degree $k$, with $k\leq d-2$.
Specializing to degree-one foliations on smooth complete intersections with $c_{1}(TX)\geq 2$, we conclude that either the foliation is defined by a closed rational 1-form without codimension-one zeros, or it admits a codimension-two subfoliation of degree zero (see Corollary 4.9). In the latter case, we prove that the foliation is the pullback of a degree-one foliation on the projective plane, and is therefore again defined by a closed rational 1-form without codimension-one zeros. This completes the proof of Theorem A.
Finally, we apply the same techniques to analyze the structure of degree-two foliations on smooth quadrics (see Corollary 4.10 and Theorem 4.11), providing a partial classification for this setting and laying groundwork for future results.
This article is organized as follows. In Section 2, we introduce the main definitions and background on holomorphic foliations, and describe the space of codimension-one foliations on projective varieties. Section 3 is devoted to the study of extensions and restrictions of foliations. In Section 4, we develop the tangential foliation by analyzing deformations of rational curves along the leaves, and use this to prove structural results for codimension-one foliations of degrees one and two on complete intersections. In Section 5, we apply the previous results to complete the classification of degree-one foliations on complete intersections.
2. Holomorphic Foliations
Let $X$ be a smooth projective complex manifold. A distribution $\mathcal{F}$ on $X$ is given by a saturated subsheaf ${T_{\mathcal{F}}}\subset TX$, often called the tangent sheaf of $\mathcal{F}$. Here, by saturated we mean that the quotient $TX/{T_{\mathcal{F}}}$ is a torsion-free sheaf. The singular locus of $\mathcal{F}$, denoted by $\operatorname{Sing}(\mathcal{F})$, is the locus where the quotient $TX/{T_{\mathcal{F}}}$ is not locally free. Since $TX/{T_{\mathcal{F}}}$ is torsion-free, it follows that $\operatorname{codim}\operatorname{Sing}(\mathcal{F})\geq 2$. We define the normal sheaf of $\mathcal{F}$ as ${N_{\mathcal{F}}}=(TX/{T_{\mathcal{F}}})^{\vee\vee}$.
The tangent sheaf ${T_{\mathcal{F}}}$ is involutive if it is closed under the Lie bracket of vector fields; in this case, $\mathcal{F}$ defines a foliation, i.e., through every point $x\in X\setminus\operatorname{Sing}(\mathcal{F})$ passes an immersed analytic submanifold $\mathcal{L}$ of dimension $\operatorname{rk}{T_{\mathcal{F}}}$ such that $T_{x}\mathcal{L}=T\mathcal{F}_{x}$; these are the leaves of the foliation.
The map ${TX}\to{N_{\mathcal{F}}}$ is defined by contraction with a twisted $p$-form $\omega_{\mathcal{F}}\in H^{0}(X,\Omega^{p}_{X}\otimes\det({N_{\mathcal{F}}}))$, where $p=\operatorname{rk}{N_{\mathcal{F}}}$. Conversely, given a line bundle $L\in\operatorname{Pic}(X)$, $\omega\in H^{0}(X,\Omega^{p}_{X}\otimes L)$ defines a distribution if and only if it is locally decomposable off the singular set, LDS for short, see [11]. This condition means that for every point $x\in X\setminus\operatorname{Sing}(\omega)$ there exists an open neighborhood $U\ni x$ and $1$-forms $\alpha_{1},\dots,\alpha_{p}$ such that
$\omega|_{U}=\alpha_{1}\wedge\cdots\wedge\alpha_{p}.$
Moreover, $\omega$ is integrable (i.e., defines a foliation) if it is LDS and satisfies, for every $j=1,\dots,p$,
$d\alpha_{j}\wedge\alpha_{1}\wedge\cdots\wedge\alpha_{p}=0.$
The local decomposability and integrability condition determine algebraic equations on ${\mathbb{P}}H^{0}(X,\Omega^{p}_{X}\otimes L)$ defining the space of foliations
${\sf Fol}^{p}(X,L)\subset{\mathbb{P}}H^{0}(X,\Omega^{p}_{X}\otimes L).$
The main objective of our work is to describe the irreducible components of this space for $p=1$, $X\subset{\mathbb{P}^{n+1}}$ a smooth complete intersection and specified $L$. Any line bundle on ${\mathbb{P}^{n}}$ is isomorphic to ${\mathcal{O}}_{\mathbb{P}^{n}}(k)$ for some $k\in{\mathbb{Z}}$. A distribution has degree $d$ if it is defined by $\omega\in H^{0}(\Omega_{\mathbb{P}^{n}}^{p}(d+p+1))$. This follows Lins Neto’s convention of the degree of a foliation. In general, if $X$ is a smooth projective variety such that $\operatorname{Pic}(X)={\mathbb{Z}}$, we define the degree of a foliation $\mathcal{F}$ determined by a section in $H^{0}(X,\Omega_{X}^{p}(d+p+1))$ as $\deg(\mathcal{F}):=d$. In this case, for simplicity, we will denote ${\sf Fol}(X,d):={\sf Fol}^{1}(X,\mathcal{O}_{X}(d+2))$. We are particularly interested in smooth complete intersections.
Smooth complete intersections of dimension at least three in ${\mathbb{P}^{n}}$ have Picard group isomorphic to ${\mathbb{Z}}$. This is due, essentially, to the Lefschetz Hyperplane Theorem and the following fact.
Lemma 2.1.
Let $X\subset{\mathbb{P}^{N}}$ be a smooth complete intersection of codimension $k$. Then there exist hypersurfaces $D_{1},\dots,D_{k}\subset{\mathbb{P}^{N}}$ such that $X_{i}=D_{1}\cap\dots\cap D_{i}$ is smooth for $i=1,\dots,k$ and $X=X_{k}$.
Proof.
Assume $X=V(f_{1},\dots,f_{k})$ with $\deg(f_{1})\geq\deg(f_{2})\geq\cdots\geq\deg(f_{k})$; denote $d_{i}=\deg(f_{i})$. For each $1\leq i<j\leq k$, let $h_{ij}$ be a polynomial of degree $d_{i}-d_{j}$, to be chosen generically. Then define
$g_{i}=f_{i}+\sum_{j=i+1}^{k}h_{ij}f_{j}\quad\text{and}\quad X_{i}=V(g_{1},\dots,g_{i});$
clearly $X=X_{k}$. We need to show that $X_{i}$ is smooth for a generic choice of $h_{ij}$. We proceed by induction. For the sake of the argument, define $X_{0}={\mathbb{P}^{n}}$. Suppose $X_{i-1}$ is smooth, for some $i\geq 1$.
Note that $X_{i}=X_{i-1}\cap V(g_{i})$ is an element of the linear system $\Lambda$ spanned by $f_{i}$ and $f_{j}h_{ij}$, for $h_{ij}$ varying in a basis of $H^{0}({\mathcal{O}}_{X_{i-1}}(d_{i}-d_{j}))$. By Bertini’s Theorem, $X_{i}$ can be chosen smooth away from the base locus, which is precisely $X$. Now, if $X_{i}$ were singular at some point $x\in X$, then $X=X_{i}\cap V(f_{i+1},\dots,f_{k})$ would also be singular at $x$, which is absurd. ∎
Remark 2.1.
In [1, 14, 15] the authors study distributions and foliations on smooth complete intersections satisfying an extra hypothesis: $X=V(f_{1},\dots,f_{k})$ with $\deg(f_{1})\leq\dots\leq\deg(f_{k})$ such that $X_{i}=V(f_{1},\dots,f_{i})$ is smooth for every $i$. There are smooth complete intersections without this property; for instance, $X=V(f_{1},f_{2})$ where $\deg(f_{1})<\deg(f_{2})$ and $V(f_{1})$ is singular. However, one can verify that the arguments in these papers depend uniquely on the existence of a chain of smooth complete intersection $X_{1}\supset\cdots\supset X_{k}=X$, regardless of the degrees. Therefore, in light of Lemma 2.1, their results apply to smooth complete intersections in general.
Remark 2.2.
Note further that a smooth complete intersection only admits distributions and foliations of nonnegative degree. Indeed, $H^{0}(X,\Omega_{X}^{p}(d+p+1))=0$ for $d<0$ and $X$ a smooth complete intersection, see [1, Lemma 5.17].
2.1. Algebraically integrable foliations
A foliation on $X$ is algebraically integrable if its general leaf is an algebraic subvariety. In this case, there exists a dominant rational map $\pi\colon X\dashrightarrow Y$ onto a projective variety $Y$ such that the closure of a leaf is an irreducible component of a fiber of $\pi$, see [2, §3].
Proposition 2.2.
Let $X$ be a complex projective manifold of dimension $n\geq 3$ with $\operatorname{Pic}(X)\cong{\mathbb{Z}}$. Let $\mathcal{F}$ be a codimension-one foliation on $X$. Suppose that $\mathcal{F}$ is algebraically integrable, then either:
  1. (1)
    $\mathcal{F}$ admits a rational first integral with only irreducible fibers; or
  2. (2)
    there is a codimension-two subfoliation $\mathcal{G}\subset\mathcal{F}$ such that $c_{1}({N_{\mathcal{F}}})=c_{1}({N_{\mathcal{G}}})+\Delta$, with $\Delta\geq 0$.
Proof.
This follows directly from the arguments in [20, Lemma 3.1 and Theorem 3.2]. ∎
Corollary 2.3.
Let $X\subset{\mathbb{P}^{N}}$ be a smooth complete intersection of dimension $n\geq 3$ and $\mathcal{F}$ be a codimension-one singular holomorphic foliation on $X$ of degree $d$. If $\mathcal{F}$ is algebraically integrable, then
  1. (1)
    $\mathcal{F}$ is defined by a rational $1$-form without zeros in codimension-one; or
  2. (2)
    $\mathcal{F}$ admits a codimension-two subfoliation $\mathcal{G}$ of degree $k$ such that $k\leq d-1$.
Proof.
By Proposition 2.2, either $\mathcal{F}$ admits a rational first integral $\rho\colon X\dashrightarrow{\mathbb{P}^{1}}$ with only irreducible fibers or it admits a subfoliation $\mathcal{G}$. Suppose we are in the first case. By Lefschetz Hyperplane Theorem, $X$ is simply connected. Then the argument follows the proof of [20, Proposition 3.5]. By [20, Theorem 3.3], $\rho$ admits at most $2$ multiple fibers, which we may assume to be over 00 and $\infty$, and write $\rho^{-1}(0)=pH_{0}$ and $\rho^{-1}(\infty)=qH_{\infty}$. Then $\rho^{*}(\frac{dx}{x}-\frac{dy}{y})$ is a logarithmic $1$-form defining $\mathcal{F}$ with polar divisor $H_{0}+H_{\infty}$ and empty zero divisor. ∎
Notice that the proof of corollary holds for any simply connected manifold $X$ with Picard number $1$. For our purposes, we will only need $X$ to be a smooth complete intersection, especially hypersurfaces.
2.2. Irreducible components of the space of foliations
A degree-zero codimension-one foliation on ${\mathbb{P}^{n}}$ is determined by a twisted $1$-form $\omega\in H^{0}({\mathbb{P}^{n}},\Omega_{\mathbb{P}^{n}}^{1}(2))$. Any of them can be written as
$\omega=L_{1}dL_{2}-L_{2}dL_{1},$
where $L_{1}$ and $L_{2}$ are degree one homogeneous polynomials, and thus they are all conjugate to each other (see [13, Proposition 3.1]). Therefore, ${\sf Fol}({\mathbb{P}^{n}},0)$ is isomorphic to the Grassmannian of lines in ${\mathbb{P}^{n}}$.
For codimension-one foliations of degree one, Jouanoulou in [16] proved that ${\sf Fol}({\mathbb{P}^{n}},1)$ has two irreducible components. Both components parameterize logarithmic foliations, which we now describe in more generality.
Let $X$ be any smooth projective variety such that $\dim(X)\geq 3$, $\operatorname{Pic}(X)={\mathbb{Z}}$ and $H^{1}(X,{\mathbb{C}})=0$. Given $d_{1},d_{2},\dots,d_{k}$ positive integers and $\lambda_{1},\lambda_{2},\dots,\lambda_{k}$ complex numbers, not all zero, such that $\sum\lambda_{i}\cdot d_{i}=0$, we can take sections $f_{i}\in H^{0}(X,{\mathcal{O}}_{X}(d_{i}))$ to form global sections
(2.1) $\omega=\left(\prod^{k}_{i=1}f_{i}\right)\sum^{k}_{i=1}\lambda_{i}\dfrac{df_{i}}{f_{i}}\in H^{0}(X,\Omega^{1}_{X}(\sum d_{i})).$
Then $\omega$ is integrable and, if its singular locus has codimension greater than one, it determines a foliation known as a logarithmic foliation. Varying $\lambda_{i}$ and $f_{i}$, $i=1,\dots,k$, the forms $\omega$ sweep an irreducible algebraic subvariety of ${\sf Fol}(X,d_{1}+\dots+d_{k}-2)$ whose closure is denoted by $\textsf{Log}(X;d_{1},d_{2},\dots,d_{k})$. In [6], it was established that $\textsf{Log}(X;d_{1},d_{2},\dots,d_{k})$ is an irreducible component of ${\sf Fol}(X,\sum d_{i}-2)$. It is called the logarithmic component. If $k=2$, it is also called a rational component. These varieties can also be defined by the closure of the rational map
$\begin{array}[]{crcl}\Phi_{X}\colon&\Sigma\times\displaystyle\prod^{k}_{i=1}{\mathbb{P}}H^{0}(X,{\mathcal{O}}_{X}(d_{i}))&\dashrightarrow&\mathbb{P}H^{0}(X,\Omega^{1}_{X}(\sum d_{i}))\\ &((\lambda_{1}:\dots:\lambda_{k}),f_{1},\dots,f_{k})&\mapsto&\left(\displaystyle\prod^{k}_{i=1}f_{i}\right)\left(\displaystyle\sum^{k}_{i=1}\lambda_{i}\dfrac{df_{i}}{f_{i}}\right)\end{array},$
where $\Sigma\subset{\mathbb{P}^{k-1}}$ is the hyperplane defined by $\sum\lambda_{i}\cdot d_{i}=0$, if its domain is not empty.
For ${\sf Fol}({\mathbb{P}^{n}},1)$, the two irreducible components are the rational component $\textsf{Log}({\mathbb{P}^{n}};2,1)$ and the logarithmic component $\textsf{Log}({\mathbb{P}^{n}};1,1,1)$.
We can also describe $\textsf{Log}({\mathbb{P}^{n}};1,1,1)$ as linear pullbacks of degree-one foliations on ${\mathbb{P}^{2}}$. In general, consider $\textsf{LPB}(d)$ the closure of the image of the map
$\begin{array}[]{rcl}\Theta\colon\mathbb{L}_{2}\times{\sf Fol}({\mathbb{P}^{2}},d)&\longrightarrow&{\sf Fol}({\mathbb{P}^{n}},d)\\ (\pi,\mathcal{G})&\longmapsto&\pi^{*}\mathcal{G}\end{array},$
where $\mathbb{L}_{2}$ is the set of linear projections of ${\mathbb{P}^{n}}$ to ${\mathbb{P}^{2}}$. Then, in [8], it was demonstrated that $\textsf{LPB}(d)$ is an irreducible component of ${\sf Fol}({\mathbb{P}^{n}},d)$, and it is called the Linear Pull-Back component. It is an elementary exercise to show that $\textsf{LPB}(1)=\textsf{Log}({\mathbb{P}^{n}};1,1,1)$.
For degree two, Cerveau and Lins Neto proved in [7] that ${\sf Fol}({\mathbb{P}^{n}},2)$ has six irreducible components: a pull-back Linear component $\textsf{LPB}(2)$; the rational components $\textsf{Log}({\mathbb{P}^{n}};3,1)$ and $\textsf{Log}({\mathbb{P}^{n}};2,2)$; the logarithmic components $\textsf{Log}({\mathbb{P}^{n}};2,1,1)$ and $\textsf{Log}({\mathbb{P}^{n}};1,1,1,1)$; and another irredutible component called exeptional component, which is determined by the closure of the map
$\begin{array}[]{rcl}\Psi\colon\mathbb{L}_{3}&\rightarrow&{\sf Fol}({\mathbb{P}^{n}},2)\\ \pi&\mapsto&\pi^{*}\mathcal{E}\end{array},$
where $\mathcal{E}$ is the exceptional foliation determined by the section
$\beta=\frac{1}{x_{0}}(2fdg-3gdf)\in H^{0}({\mathbb{P}^{3}},\Omega^{1}_{{\mathbb{P}^{3}}}(4)).$
where $f=2x_{0}x_{2}-x_{1}^{2}$ and $g=3x_{0}^{2}x_{3}-3x_{0}x_{1}x_{2}+x_{1}^{3}$. In [10], the irreducible components of ${\sf Fol}({\mathbb{P}^{n}},3)$ were partially classified; there are at least 24 irreducible components.
Besides the projective space, there is some literature on irreducible components of the space of foliations on, for instance, Fano three-folds [20], homogeneous spaces [5, 19], and hypersurfaces [1, 15].
In the case of smooth projective hypersurfaces $X$ in ${\mathbb{P}^{n+1}}$, $n\geq 3$, if
$\deg(X)>2\deg(\mathcal{F})+1,$
then the restriction of foliations from the ambient ${\mathbb{P}^{n+1}}$ defines an isomorphism between ${\sf Fol}(X,d)$ and ${\sf Fol}({\mathbb{P}^{n+1}},d)$, see [15, Theorem 3.6]. Therefore, the previous classifications for foliations on projective spaces are valid for these varieties. The problem is then to study low-degree hypersurfaces; for degree-one foliations, we need to analyze hypersurfaces of degree at most three.
In [20], the authors proved that for $X\subset{\mathbb{P}^{4}}$ a smooth quadric hypersurface, ${\sf Fol}(X,1)$ has three irreducible components; only two are restrictions from foliations on ${\mathbb{P}^{4}}$. In this work, we prove that, for hypersurfaces, this is the only case where ${\sf Fol}(X,1)$ has more than two irreducible components.
2.3. Degree-zero distributions on the projective space
In [1], the authors classify distributions of codimension one and degree zero on ${\mathbb{P}^{n}}$. There is a stratification according to the class of the $1$-form defining the distribution. The argument boils down to the classification of skew-symmetric matrices and the nesting of secant varieties to the Grassmannian of lines in ${\mathbb{P}}(\textstyle\bigwedge^{2}{\mathbb{C}}^{n+1})$. In higher codimension, the situation is surprisingly simpler.
Proposition 2.4.
Let $\mathcal{D}$ be a degree-zero distribution of codimension $q\geq 2$ on ${\mathbb{P}^{n}}$. Then $\mathcal{D}$ is integrable.
Proof.
Let $\omega\in H^{0}({\mathbb{P}^{n}},\Omega_{\mathbb{P}^{n}}^{q}(q+1))$ a $q$-form defining $\mathcal{D}$. By [11, Theorem A], there exists a choice of homogeneous coordinates such that either $\omega=\alpha\wedge dx_{0}\wedge\cdots\wedge dx_{q-2}$ for some $\alpha$ a $1$-form with linear coefficients, or $\omega=i_{v}dx_{0}\wedge\cdots\wedge dx_{q}$, for some linear vector field $v$. Imposing $i_{\rm rad}\omega=0$, for ${\rm rad}=\sum_{j}x_{j}\frac{\partial}{\partial x_{j}}$ the radial vector field, we must have $\omega=i_{\rm rad}dx_{0}\wedge\cdots\wedge dx_{q}$, up to a scalar multiple. Therefore, $\mathcal{D}$ is integrable. ∎
3. Extensions and restrictions of foliations on complete intersections
Let $U_{k}$ be the open subset of twisted $q$-forms in $H^{0}({\mathbb{P}^{n}},\Omega_{\mathbb{P}^{n}}^{q}(k))$ whose singular set has codimension greater than or equal to two and let $X\subset{\mathbb{P}^{n}}$ be a smooth complete intersection.
Definition 3.1.
We say that $\omega\in U_{k}$ is transversal in codimension one to $X$ if $i^{*}\omega$ has a singular set of codimension greater than or equal to two, where $i\colon X\rightarrow{\mathbb{P}^{n}}$ is the natural inclusion. A singular holomorphic foliation $\mathcal{F}$ of codimension $q$ is transversal in codimension one to $X$, if a twisted $q$-form that determines it is transversal in codimension-one to $X$.
If a foliation $\mathcal{F}$ is transversal in codimension one to $X$, the restriction to $X$ of its twisted $q$-form naturally satisfies the decomposability and integrability hypotheses and, therefore, defines a foliation $\mathcal{G}$ on $X$.
Definition 3.2.
The foliation $\mathcal{G}$ will be called the restriction of $\mathcal{F}$ to $X$. We will denote ${\left.\kern-1.2pt\mathcal{F}\vphantom{\big{|}}\right|_{X}}:=\mathcal{G}$. If a foliation $\mathcal{G}$ on a smooth complete intersection $X\subset{\mathbb{P}^{n}}$ is the restriction of a foliation $\mathcal{F}$ in ${\mathbb{P}^{n}}$ to $X$, we say that $\mathcal{F}$ is an extension for $\mathcal{G}$.
3.1. Codimension-one foliations
Let $X\subset{\mathbb{P}^{n}}$ be a smooth complete intersection of dimension $r\geq 3$. The restriction maps
$\operatorname{rest}_{0}\colon H^{0}({\mathbb{P}^{n}},{\mathcal{O}}_{\mathbb{P}^{n}}(k))\rightarrow H^{0}(X,{\mathcal{O}}_{X}(k))$
are surjective, for all $k\in\mathbb{Z}$. Thus, if $\mathcal{F}$ is an algebraically integrable foliation on $X$ defined by a pencil of hypersurfaces sections $f,g$, where $f\in H^{0}(X,{\mathcal{O}}_{X}(\deg(f))$ and $g\in H^{0}(X,{\mathcal{O}}_{X}(\deg(g))$. Then, the 1-form $\deg(g)gdf-\deg(f)fdg$ determines $\mathcal{F}$. Therefore, we can find coprime polynomials $F,G$ such that $\operatorname{rest}_{0}(F)=f$ and $\operatorname{rest}_{0}(G)=g$, and the foliation restriction generated by $\deg(g)GdF-\deg(f)FdG$ is the extension of $\mathcal{F}$.
In general, for logarithmic forms as in (2.1), we have the commutative diagram
[…]
which projects all logarithmic forms of ${\mathbb{P}^{n}}$ onto $X$ surjectively, where $i\colon X\rightarrow{\mathbb{P}^{n}}$ is the natural inclusion map. Therefore, all foliations generated by logarithmic forms extend.
Proposition 3.1.
Let $X$ be a smooth complete intersection. Then, every codimension-one foliation on any logarithmic component $\textsf{Log}(X,d_{1},d_{2},\dots,d_{k})$ extends.
Proof.
As seen above, if $\mathcal{F}$ is generated by a logarithmic form, it has an extension. Furthermore, the logarithmic and rational components can be determined by the closure of the image of a rational map as (3.1).
Suppose $\mathcal{F}$ is in a logarithmic component $\textsf{Log}(X,d_{1},d_{2},\dots,d_{k})\subset{\sf Fol}(X,\sum d_{i}-2)$ and is an adherence value of the map $\Phi$. In this case, there exists a sequence of foliations $\{\mathcal{F}_{i}\}_{i=0}^{\infty}$ in the logarithmic component such that $\mathcal{F}_{i}\rightarrow\mathcal{F}$ and each $\mathcal{F}_{i}$ is defined by a logarithmic form. Since each foliation $\mathcal{F}_{i}$ extends, there exists a sequence of foliations $\{\mathcal{G}_{i}\}_{i=0}^{\infty}$ in the logarithmic component $\textsf{Log}({\mathbb{P}^{n}},d_{1},d_{2},\dots,d_{k})$, where ${\left.\kern-1.2pt\mathcal{G}_{i}\vphantom{\big{|}}\right|_{X}}=\mathcal{F}_{i}$ and each $\mathcal{G}_{i}$ is determined by a logarithmic form $\omega_{i}$. Since the integrability condition determines a proper closed set in the class space $\mathbb{P}H^{0}({\mathbb{P}^{n}},\Omega_{\mathbb{P}^{n}}^{1}(k)),$ the set of points adhering to the sequence $([\omega_{i}])_{i=0}^{\infty}$ is nonempty and each point in this set is integrable. Thus, given $[\omega]$ as one of these adherent points, its restriction to $X$ determines $\mathcal{F}$ by definition. Furthermore, we have that $\operatorname{codim}(\omega)\geq 2$, otherwise, $\operatorname{codim}(\operatorname{rest}(\omega))\leq 1$ and $\operatorname{rest}(\omega)$ would not determine $\mathcal{F}$. Therefore, $\omega$ determines a logarithmic foliation $\mathcal{G}$ on ${\mathbb{P}^{n}}$ whose restriction is $\mathcal{F}$. ∎
3.2. Codimension two foliations
For codimension-two foliations of degree zero, we have the following results regarding extensions.
Theorem 3.2.
Let $X$ be a smooth complete intersection in ${\mathbb{P}^{n+s}}$ of type $(d_{1},\ldots,d_{s})$ and dimension $n\geq 4$. If $d_{j}\geq 3$ for $j=1,\dots,s$, then any codimension-two foliation on $X$ of degree zero has a unique extension to ${\mathbb{P}^{n+s}}$.
Proof.
Let $\mathcal{F}$ be a codimension-two foliation on $X$ of degree zero. In particular, $\mathcal{F}$ is a codimension-two distribution on $X$ of degree zero. Applying [14, Corolário 2.3.3], we see that $\mathcal{F}$ has a unique extension as a codimension-two distribution to ${\mathbb{P}^{n+s}}$ of degree zero. Finally, by Proposition 2.4, we conclude that the extension of $\mathcal{F}$ is integrable. ∎
Corollary 3.3.
Let $X$ be a smooth hypersurface of ${\mathbb{P}^{n+1}}$, $n\geq 4$. If $\mathcal{F}$ is a codimension-two holomorphic foliation on $X$ of degree zero then $\mathcal{F}$ extends on ${\mathbb{P}^{n+1}}$.
Proof.
It follows by Theorem 3.2, if $\deg(X)\geq 3$ and, for quadrics, by [3, Proposition 3.18]. ∎
4. Structure Theorem for Foliations on Complete intersections
In this section, we prove our main technical result, Theorem 4.8. Then, we derive the structure of degree-one foliations on smooth complete intersections and degree-two foliations on quadric hypersurfaces. We start with some preparation.
4.1. Tangential foliations
A smooth projective variety $X$ is called uniruled if, for every general point $x\in X$, there exists a non-constant morphism $f\colon{\mathbb{P}^{1}}\to X$ such that the rational curve $C=f({\mathbb{P}^{1}})$ passes through $x$ and $f^{*}{TX}$ is generated by global sections; such morphisms are called free. The scheme $\operatorname{Mor}({\mathbb{P}^{1}},X)$ parametrizing morphisms from ${\mathbb{P}^{1}}$ to $X$ is smooth around a free morphism $f$ and its tangent space is naturally isomorphic to $H^{0}(X,f^{*}{TX})$. We refer to [12, 18] for further details. A particular result we will need is the following.
Lemma 4.1.
Let $f\colon{\mathbb{P}^{1}}\to X$ be a free morphism and $Z\subset X$ be a closed subvariety of codimension at least two. Then, $f^{\prime}({\mathbb{P}^{1}})\cap Z=\emptyset$ for a general deformation $f^{\prime}$ of $f$.
Proof.
This is a particular case of [18, II, Proposition 3.7]. ∎
Lemma 4.2.
Let $X$ be a smooth complex quasi-projective variety. There exists a non-empty subset $X^{\rm free}$ of $X$ which is the intersection of countably many dense open subsets of $X$, such that any rational curve on $X$ whose image meets $X^{\rm free}$ is free.
Proof.
See [12, Proposition 4.14] for proof of the existence of $X^{\rm free}$ over any algebraically closed field of characteristic zero. If the base field is uncountable, then $X^{\rm free}\neq\emptyset$. ∎
Given a foliation $\mathcal{F}$ on $X$ and an irreducible component $M\subset\operatorname{Mor}({\mathbb{P}^{1}},X)$ containing a free morphism. One defines the tangential foliation of $\mathcal{F}$ on $M$, denoted by $\mathcal{F}_{\mathrm{tang}}$, setting for a general free morphism $f:{\mathbb{P}^{1}}\dashrightarrow X$ the tangent space of $\mathcal{F}_{\mathrm{tang}}$ at $f$ equal to $H^{0}({\mathbb{P}^{1}},f^{*}T_{\mathcal{F}})$, via the identification of the tangent space of $\mathrm{Mor}(\mathbb{P}^{1},X)$ at the morphism $f$ with $H^{0}({\mathbb{P}^{1}},f^{*}T_{X})$ and the inclusion $H^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})\subset H^{0}({\mathbb{P}^{1}},f^{*}{TX})$. More precisely, the tangential foliation is defined as follows. Consider the evaluation map
$\mathrm{ev}:M\times{\mathbb{P}^{1}}\longrightarrow X$
$([f],x)\mapsto f(x),$
and the natural projection $\pi:M\times{\mathbb{P}^{1}}\longrightarrow M$. We have an isomorphism $T_{M}\cong\pi_{*}\mathrm{ev}^{*}T_{X}$, and so the inclusion $T_{\mathcal{F}}\subset T_{X}$ induces a morphism $i:\pi_{*}\mathrm{ev}^{*}T_{\mathcal{F}}\hookrightarrow T_{M}$. We define the tangential foliation of $\mathcal{F}$ on $M$ by the saturation of the image of this morphism $i$ in $T_{M}$. In [21], the authors show how the geometry of $\mathcal{F}_{\rm tang}$ impacts the geometry of $\mathcal{F}$. We recall below two of their results.
Proposition 4.3.
[21, Proposition 4.10] Let $X$ be a uniruled smooth projective variety and let $\mathcal{F}$ be a codimension one foliation on $X$. If all leaves of $\mathcal{F}_{\rm tang}$ are algebraic, then there exists an algebraically integrable foliation $\mathcal{G}$ contained in $\mathcal{F}$ such that $H^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})=H^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{G}}})$ for any sufficiently general $f\in M$. In particular, $\mathcal{F}$ is the pullback under a rational map of a foliation on a lower-dimensional manifold.
Theorem 4.4.
[21, Theorem C] Suppose that $X$ is simply connected, $\operatorname{codim}\mathcal{F}=1$, the general leaf of $\mathcal{F}_{\rm tang}$ is not algebraic, and a general $f\in M$ nontrivially and transversely all the algebraic hypersurfaces invariant by $\mathcal{F}$. Then $\mathcal{F}$ is defined by a closed rational $1$-form without divisorial components in its zero set.
Remark 4.1.
In Proposition 4.3, we observe that $f^{*}{T_{\mathcal{G}}}$ contains the nonnegative part of $f^{*}{T_{\mathcal{F}}}$. Let $a_{1}\geq\dots\geq a_{r}$ and $b_{1}\geq\dots\geq b_{s}$ be integers such that $f^{*}{T_{\mathcal{F}}}=\bigoplus_{i=1}^{r}{\mathcal{O}}_{\mathbb{P}^{1}}(a_{i})$ and $f^{*}{T_{\mathcal{G}}}=\bigoplus_{i=1}^{s}{\mathcal{O}}_{\mathbb{P}^{1}}(b_{i})$. The inclusion $f^{*}{T_{\mathcal{G}}}\subset f^{*}{T_{\mathcal{F}}}$ implies that $a_{i}\geq b_{i}$ for $i=1,\dots,s$. Let $l=\max\{\,j\mid a_{j}\geq 0\,\}$ and $m=\max\{\,j\mid b_{j}\geq 0\,\}$. The equality $H^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})=H^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{G}}})$ implies that $m=l$, and $a_{i}=b_{i}$ for $i\leq m$.
We will often consider rational curves that are not tangent to the foliation. When $\operatorname{Pic}(X)\cong{\mathbb{Z}}$, this is possible due to the following result, which was communicated to us by Jorge Vitório Pereira.
Lemma 4.5.
Let $X$ be a smooth uniruled complex projective variety such that $\operatorname{Pic}(X)\cong{\mathbb{Z}}$. If $f\colon{\mathbb{P}^{1}}\to X$ is a free morphism and $\mathcal{F}$ is a foliation on $X$, then a general deformation $f^{\prime}$ of $f$ is not $\mathcal{F}$-invariant.
Proof.
Due to Lemma 4.1, we may choose $f^{\prime}$ so that $f^{\prime}({\mathbb{P}^{1}})$ does not intersect $\operatorname{Sing}(\mathcal{F})$. Aiming for a contradiction, suppose $f^{\prime}$ is tangent to $\mathcal{F}$. Recall Bott’s partial connection $\nabla^{B}\colon{T_{\mathcal{F}}}\otimes{N_{\mathcal{F}}}\to{N_{\mathcal{F}}}$ that is defined as follows. Let $v\in{T_{\mathcal{F}}}$ and $\overline{w}\in{N_{\mathcal{F}}}$ be local sections, and $w\in{TX}$ a lift of $w$. Then $\nabla^{B}_{v}(\overline{w})=\overline{[v,w]}$. Since $f^{\prime}$ is tangent to $\mathcal{F}$, we get that $f^{\prime*}\nabla^{B}$ is a holomorphic connection on the vector bundle $f^{\prime*}{N_{\mathcal{F}}}$ on ${\mathbb{P}^{1}}$. By Weil’s Theorem [4, Theorem 10], $f^{\prime*}{N_{\mathcal{F}}}$ must be trivial. Using that $\operatorname{Pic}(X)\cong{\mathbb{Z}}$, we conclude that $c_{1}({N_{\mathcal{F}}})=0$, hence $\mathcal{F}$ is given by a global holomorphic $p$-form $\omega\in H^{0}(X,\Omega^{p}_{X})$, $p=\operatorname{codim}\mathcal{F}$. On the other hand, $X$ is uniruled with $\operatorname{Pic}(X)\cong{\mathbb{Z}}$ implies that $X$ is Fano, hence $h^{0}(X,\Omega^{p}_{X})=h^{0,p}(X)=0$, which is a contradiction. ∎
Next, we turn to a particular case of uniruled varieties: varieties covered by lines. Let $L\in\operatorname{Pic}(X)$ be an ample line bundle. A morphism $f\colon{\mathbb{P}^{1}}\to X$ is called a line (with respect to $L$) if
$\int_{{\mathbb{P}^{1}}}f^{*}c_{1}(L)=1.$
The usual setting is when $X\subset{\mathbb{P}^{n}}$ and $L={\mathcal{O}}_{X}(1)$, meaning that $f({\mathbb{P}^{1}})$ is a genuine line of ${\mathbb{P}^{n}}$ contained in $X$. A variety $X$ is said to be covered by lines if there exists a line passing through a general point of $X$. It follows from Lemma 4.2 that every line passing through a point in $X^{\mathrm{free}}$ must be free. The following corollary is an immediate consequence of Proposition 4.5.
Corollary 4.6.
Let $X$ be a smooth complex projective variety covered by lines such that $\operatorname{Pic}(X)\cong{\mathbb{Z}}$. Let $\mathcal{F}$ be a foliation on $X$, then a general free line is not $\mathcal{F}$-invariant.
Suppose $X\subset{\mathbb{P}^{N}}$ is a complete intersection given by equations of degrees $d_{1},\dots,d_{s}$. If $d_{1}+\cdots+d_{s}\leq N-1$ then $X$ is covered by lines, see [12, Proposition 2.13]. Then, the previous corollary applies.
Proposition 4.7.
Let $X\subset{\mathbb{P}^{N}}$ be a smooth complex complete intersection of dimension at least three defined by equations of degrees $d_{1},\dots,d_{s}$ such that $d_{1}+\cdots+d_{s}\leq N-1$. For any foliation $\mathcal{F}$ on $X$, a general line is not $\mathcal{F}$-invariant.
4.2. Foliations of degrees one and two
Following the discussion above, we will prove a structure theorem for foliations of small degrees on smooth projective varieties covered by lines. Then, we will apply it to smooth complete intersections and, in particular, to smooth hypersurfaces. The result below is inspired by the structure theorem for degree-three foliations on ${\mathbb{P}^{3}}$ by da Costa, Lizarbe, and Pereira, see [10, Theorem A].
Theorem 4.8.
Let $X\subset{\mathbb{P}^{N}}$ be a smooth projective variety covered by lines of dimension $n\geq 3$ with $\operatorname{Pic}(X)=\mathbb{Z}$. If $\mathcal{F}$ is a codimension-one singular holomorphic foliation on $X$ of degree $d$ such that either $d=1$ or
(4.1) $c_{1}(TX)\geq n+d-2,$
then one of the following holds:
  1. (1)
    $\mathcal{F}$ is defined by a closed rational 1-form without Codimension one zeros; or
  2. (2)
    $\mathcal{F}$ is algebraically integrable; or
  3. (3)
    $\mathcal{F}$ admits an algebraically integrable subfoliation of codimension 2, and degree $k$ such that $k\leq d-2$.
Note that the inequality $c_{1}(TX)\geq n+d-2$ for $d\geq 2$ only holds for the quadric hypersurface or ${\mathbb{P}^{n}}$, due to the Kobayashi-Ochiai Theorem [17].
Proof.
By the adjunction formula, we have
$c_{1}({T_{\mathcal{F}}})=c_{1}(TX)-(d+2).$
Let $f\colon{\mathbb{P}^{1}}\to{\mathbb{P}^{N}}$ be the parameterization of a free line on ${\mathbb{P}^{N}}$ contained in $X$. Due to Proposition 4.7, we can assume that $f$ is generically transverse to $\mathcal{F}$. Consider the decomposition $f^{*}{T_{\mathcal{F}}}=\bigoplus_{i=1}^{n-1}{\mathcal{O}}_{\mathbb{P}^{1}}(a_{i})$. Moreover,
$f^{*}TX={\mathcal{O}}_{{\mathbb{P}^{1}}}(2)\oplus{\mathcal{O}}_{{\mathbb{P}^{1}}}(1)^{a}\oplus{\mathcal{O}}_{{\mathbb{P}^{1}}}^{b}$
where $a=c_{1}(TX)-2\geq 0$ and $b=n-c_{1}(TX)+1\geq 0$ see [12, Proposition 2.13 (b)]. Since $f$ is generically transverse to $\mathcal{F}$, every summand of the decomposition of $f^{*}{T_{\mathcal{F}}}$ satisfies $a_{i}\leq 1$. From
(4.2) $\sum_{i=1}^{n-1}a_{i}=c_{1}(TX)-(d+2)\geq-1,$
we have that $h^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})\neq 0$, hence $\mathcal{F}_{\rm tang}$ is well-defined. Therefore, we have the following possibilities for $f^{*}{T_{\mathcal{F}}}$:
  1. (1)
    $f^{*}{T_{\mathcal{F}}}$ is nonnegative in the sense that it is a direct sum of holomorphic line bundles of degree $\geq 0$.
  2. (2)
    $f^{*}{T_{\mathcal{F}}}\simeq{\mathcal{O}}_{\mathbb{P}^{1}}(1)^{r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}^{\oplus n-2-r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}(-c)$, for some nonnegative integer $r$ and $c=1$ or $c=2$.
Suppose that $\mathcal{F}$ is not algebraically integrable. We now divide the proof into the cases whether $\mathcal{F}_{\rm tang}$ is algebraically integrable or not. If $\mathcal{F}_{\rm tang}$ is not algebraically integrable, then by [21, Theorem C], $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros. If $\mathcal{F}_{\rm tang}$ is algebraically integrable, then, as argued in the proof of [21, Theorem A], there exists an algebraically integrable subfoliation $\mathcal{G}\subset\mathcal{F}$ of degree $k$ and codimension $q\geq 2$ such that $h^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{G}}})=h^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})$. It follows that $f^{*}{T_{\mathcal{F}}}\simeq{\mathcal{O}}_{\mathbb{P}^{1}}(1)^{r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}^{\oplus n-2-r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}(-c)$ and $f^{*}{T_{\mathcal{G}}}={\mathcal{O}}_{\mathbb{P}^{1}}(1)^{r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}^{\oplus n-2-r}$. Again, by the adjunction formula for $\mathcal{F}$ and $\mathcal{G}$ we have
$c_{1}({T_{\mathcal{F}}})=r-c=c_{1}(TX)-(d+2)$, and $c_{1}({T_{\mathcal{G}}})=r=c_{1}(TX)-(k+3)$.
Combining and simplifying, we get $k=d-1-c$. This proves the theorem. ∎
The previous theorem allows us to determine the structure of degree-one foliations on smooth complete intersections.
Corollary 4.9.
Let $X$ be a smooth complete intersection of ${\mathbb{P}^{N}}$ of dimension $n\geq 3$. Suppose $c_{1}(TX)\geq 2$. If $\mathcal{F}$ is a codimension-one foliation on $X$ of degree one, then
  1. (1)
    $\mathcal{F}$ is defined by a closed rational 1-form without Codimension-one zeros; or
  2. (2)
    $\mathcal{F}$ admits a subfoliation of codimension 2, and degree zero.
Proof.
Note that $X$ is covered by lines, see Proposition 4.7. By Theorem 4.8, either $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros, is algebraically integrable or has a codimension-two subfoliation of negative degree. The latter is not possible, as observed in Remark 2.2. Suppose $\mathcal{F}$ is not given by a closed rational 1-form without codimension-one zeros. Then, by Corollary 2.3, $\mathcal{F}$ must have a codimension-two subfoliation of degree zero. ∎
Corollary 4.10.
Let $\mathcal{F}$ be a codimension-one singular holomorphic foliation of degree two on a smooth quadric ${Q^{n}}\subset{\mathbb{P}^{n+1}}$, $n\geq 4$. Then
  1. (1)
    $\mathcal{F}$ is defined by a closed rational 1-form without Codimension-one zeros; or
  2. (2)
    $\mathcal{F}$ is a linear pull-back of a degree 2 foliation on ${\mathbb{P}^{2}}$; or
  3. (3)
    $\mathcal{F}$ admits a subfoliation of codimension 2, and degree one.
Moreover, foliations in item (3) are algebraically integrable.
Proof.
Suppose $\mathcal{F}$ is not defined by a closed rational 1-form without Codimension-one zeros. Combining Theorem 4.8 and Corollary 2.3, we have that $\mathcal{F}$ admits a subfoliation $\mathcal{G}$ of codimension two and degree $k\leq 1$. In case $k=0$, by [3, Proposition 3.18], $\mathcal{G}$ must be given by a linear projection to ${\mathbb{P}^{2}}$, hence $\mathcal{F}$ is a linear pullback of a foliation of degree two on ${\mathbb{P}^{2}}$. ∎
Theorem 4.11.
Let $\mathcal{F}$ be a codimension-one singular holomorphic foliation of degree two on a smooth tridimensional quadric ${Q^{3}}$. Then
  1. (1)
    $\mathcal{F}$ is defined by a closed rational 1-form without Codimension-one zeros; or
  2. (2)
    $\mathcal{F}$ is tangent to a foliation by curves of degree one; or
  3. (3)
    there exists an algebraic action of $\mathbb{C}$ or $\mathbb{C}^{*}$ tangent to $\mathcal{F}$ such that when this action has only isolated fixed points, we have ${T_{\mathcal{F}}}=\mathcal{O}_{{Q^{3}}}\oplus\mathcal{O}_{{Q^{3}}}(-1)$.
Moreover, foliations in item (2) are algebraically integrable.
Proof.
The proof is similar to the proof of Corollary 4.10. Except for the subfoliation $\mathcal{G}$ has dimension one, and [3, Proposition 3.18] does not apply. Instead, ${T_{\mathcal{G}}}={\mathcal{O}}_{Q^{3}}$ and $\mathcal{G}$ is defined by a global vector field on ${Q^{3}}$ without zeros in Codimension-one.
Let $\omega\in H^{0}({Q^{3}},\Omega_{{Q^{3}}}^{1}(3))$ be a 1-form defining $\mathcal{F}$ and define
$\mathfrak{aut}(\mathcal{F})=\{v\in H^{0}({Q^{3}},T{Q^{3}})\mid\mathcal{L}_{v}\omega\wedge\omega=0\},$
the Lie algebra of infinitesimal automorphisms of $\mathcal{F}$, here $\mathcal{L}_{v}\omega$ is the Lie derivative. Moreover, define $\mathfrak{fix}(\mathcal{F})=H^{0}({Q^{3}},T{Q^{3}})\subset\mathfrak{aut}(\mathcal{F})$. Note that $\mathcal{G}$ is defined by a vector field $v\in\mathfrak{fix}(\mathcal{F})$.
By [20, Lemma 4.3], either $\mathfrak{fix}(\mathcal{F})\neq\mathfrak{aut}(\mathcal{F})$ and $\mathcal{F}$ is defined by a closed rational 1-form without zeros in Codimension-one; or $\mathfrak{fix}(\mathcal{F})=\mathfrak{aut}(\mathcal{F})$ and there exists a non-trivial algebraic action of a one-dimensional Lie group tangent to $\mathcal{F}$; on ${Q^{3}}$ this group must be $\mathbb{C}$ or $\mathbb{C}^{*}$. Suppose we are in the latter case and assume $\mathcal{G}$ is defined by said algebraic action.
Finally, $\mathcal{G}$ has only zero-dimensional singularities, then [20, Lemma 4.2] implies that ${T_{\mathcal{F}}}={\mathcal{O}}_{Q^{3}}\oplus{\mathcal{O}}_{Q^{3}}(-1)$. ∎
5. Irreducible Components of Degree-one Foliations
In this section, we apply the structure Theorem 4.8 and its corollaries to determine the irreducible components of ${\sf Fol}(X,1)$ for $X$ either a hypersurface or smooth complete intersection of hypersurfaces of degrees at least three, improving [14, Corolário 2.2.7].
We start with a technical result.
Lemma 5.1.
Let $\pi\colon\mathbb{P}^{n+s}\dasharrow\mathbb{P}^{q}$ be a linear projection, with $n>3$ and $q<n$, and let $X\subset\mathbb{P}^{n+s}$ be a smooth complete intersection. Then the restriction $\pi|_{X}$ is a surjective map and the set of its critical points has codimension greater than one.
Proof.
First, note that $\pi|_{X}$ is surjective. In fact, as $12\leq q\leq n-$ 1, we have that the fibers of $\pi$ have dimension at least two, and the intersection of each of these fibers with $X$ is certainly not empty.
Let $\Theta\in H^{0}(\mathbb{P}^{q},\Omega^{q}_{\mathbb{P}^{q}}(q+1))$ be a $q$-twisted form in $\mathbb{P}^{q}$ without zeros; this is possible since $\Omega^{q}_{\mathbb{P}^{q}}(q+1)\cong{\mathcal{O}}_{\mathbb{P}^{q}}$. The set of critical points of $\pi|_{X}$ coincides with the zeros of $(\pi|_{X})^{*}\Theta$ which is a section of $H^{0}(X,\Omega^{q}_{X}(q+1))$. As $q<n$, we have that this section cannot vanish in codimension one because, in this case, we would produce a non-zero section in $H^{0}(X,\Omega^{q}_{X}(r))$, with $r\leq q$, but such a group is trivial. ∎
Theorem 5.2.
Let $X$ be a smooth complete intersection in ${\mathbb{P}^{n+s}}$ of type $(d_{1},\ldots,d_{s})$, with $n\geq 4$ and $c_{1}(TX)\geq 2$. If either $s=1$ ($X$ is a hypersurface) or $d_{j}\geq 3$ for $j=1,\dots,s$, then the ${\sf Fol}(X,1)$ has two irreducible components:
${\sf Fol}(X,1)=\textsf{Log}(X;1,2)\cup\textsf{Log}(X;1,1,1).$
Proof.
Let $\mathcal{F}$ be a codimension-one foliation of degree one on $X$. By Corollary 4.9, we have that $\mathcal{F}$ is either defined by a closed rational 1-form without codimension one zeros or $\mathcal{F}$ admits a subfoliation of codimension 2, and degree zero. In the first case, $\mathcal{F}$ belongs to $\textsf{Log}(X;1,2)$ or $\textsf{Log}(X;1,1,1)$, see [20, Lemma 2.5].
Suppose there exists a subfoliation $\mathcal{G}\subset\mathcal{F}$ of codimension two and degree zero. By Theorem 3.2, and Corollary 3.3 in the case of hypersurfaces, we have that $\mathcal{G}$ extends to ${\mathbb{P}^{n+s}}$ and thus $\mathcal{G}$ is defined by the restriction to $X$ of a linear projection $\pi\colon{\mathbb{P}^{n+s}}\dashrightarrow{\mathbb{P}^{2}}$. By [9, Lemma 3.1], there exists a foliation $\mathcal{H}$ on ${\mathbb{P}^{2}}$ such that $\mathcal{F}=({\left.\kern-1.2pt\pi\vphantom{\big{|}}\right|_{X}})^{*}\mathcal{H}$. Furthermore, Lemma 5.1 guarantees that ${N_{\mathcal{F}}}=({\left.\kern-1.2pt\pi\vphantom{\big{|}}\right|_{X}})^{*}{N_{\mathcal{H}}}$. Therefore, $\mathcal{H}$ is a codimension-one foliation of degree one on ${\mathbb{P}^{2}}$. The classification of foliations of degree one on ${\mathbb{P}^{2}}$ implies that $\mathcal{F}$ belongs to either $\textsf{Log}(X;1,2)$ or $\textsf{Log}(X;1,1,1)$. ∎
Note that for $X\subset{\mathbb{P}^{n+1}}$ a hypersurface, $c_{1}(TX)=n+2-\deg(X)$. Thus, Theorem 5.2 applies when $\deg(X)\leq n$. On the other hand, [15, Theorem 3.6] shows that every degree-one foliation on $X$ extends to ${\mathbb{P}^{n+1}}$ if $\deg(X)\geq 4$ and $n\geq 3$. The following is the immediate.
Corollary 5.3.
Let $X$ be a smooth hypersurface of ${\mathbb{P}^{n+1}}$, such that either $n\geq 4$ or $n=3$ and $\deg(X)\geq 4$. The irreducible components of the space of codimension-one foliations of degree one on $X$ are $\textsf{Log}(X;1,2)$ and $\textsf{Log}(X;1,1,1)$.
Corollary 5.4.
Every codimension-one foliation of degree one on a smooth hypersurface of dimension at least four extends to ${\mathbb{P}^{n+1}}$.
Proof.
This follows by Corollary 5.3 and Proposition 3.1. ∎
To complete the classification of ${\sf Fol}(X,1)$ for any smooth projective hypersurface $X$ of dimension $n\geq 3$, we only need to analyze the case of cubic hypersurfaces of dimension three. The case of a quadric threefold was described in [20, Theorem 5.2]. In addition to the logarithmic components, there is an irreducible component whose general element is given by an action of the affine group $\mathfrak{aff}(\mathbb{C})$. Furthermore, Corollary 5.3 provides us with the classification for the other hypersurfaces.
5.0.1. Foliations on a cubic threefold
Let $X\subset{\mathbb{P}^{4}}$ be a smooth cubic hypersurface and $\mathcal{F}$ a codimension-one degree-one foliation on $X$ given by $\omega_{0}\in H^{0}(X,\Omega_{X}^{1}(3))$. This form extends to (a unique!) $\omega\in H^{0}({\mathbb{P}^{4}},\Omega_{\mathbb{P}^{4}}^{1}(3))$ defining a degree-one distribution on ${\mathbb{P}^{4}}$. If $\omega$ is integrable then $\mathcal{F}$ is the restriction of a degree-one foliation on ${\mathbb{P}^{4}}$, thus $\mathcal{F}$ belongs to either $\textsf{Log}(X;2,1)$ or $\textsf{Log}(X;1,1,1)$. We will prove this is always the case.
Theorem 5.5.
A codimension-one foliation of degree one on a smooth cubic threefold $X\subset{\mathbb{P}^{4}}$ extends to a degree-one codimension-one foliation on ${\mathbb{P}^{4}}$.
Proof.
Let $\mathcal{F}$ be a codimension-one foliation of degree one on $X$, and $\omega\in H^{0}({\mathbb{P}^{4}},\Omega_{\mathbb{P}^{4}}^{1}(3))$ be such that $\omega|_{X}$ defines $\mathcal{F}$. Suppose by contradiction that $\omega$ is not integrable, that is, $\omega\wedge d\omega\neq 0$. Then there exists a polynomial vector field $v\neq 0$ such that $\omega\wedge d\omega=\iota_{{\rm rad}}\iota_{v}(dx_{0}\wedge\dots\wedge dx_{4})$, where ${\rm rad}=\sum_{j=0}^{4}x_{j}\frac{\partial}{\partial x_{j}}$ is the radial vector field. Let $f\in{\mathbb{C}}[x_{0},\dots,x_{4}]_{3}$ be a polynomial defining $X$. Since $(\omega\wedge d\omega)|_{X}=0$, we may assume $v(f)=0$, up to adding to $v$ a multiple of ${\rm rad}$. Note that, since $X$ is smooth, the partial derivatives of $f$ form a regular sequence. Hence, the Koszul complex associated with the gradient $\nabla(f)$ is exact, and $v$ must be of the form
(5.1) $v=\sum_{i,j}\frac{\partial f}{\partial x_{i}}a_{ij}\frac{\partial}{\partial x_{j}},$
where $a_{ji}=-a_{ij}\in{\mathbb{C}}$. The action of changing coordinated by $\phi\in\rm GL_{5}$ transforms $v$ as follows:
$\phi_{*}v$ $=\sum_{i,j,l}\phi_{jl}\frac{\partial f}{\partial x_{i}}\circ\phi^{-1}a_{il}\frac{\partial}{\partial x_{j}}=\sum_{i,j,l}\phi_{jl}\frac{\partial f}{\partial x_{i}}\circ\phi^{-1}a_{il}\frac{\partial}{\partial x_{j}}$
$=\sum_{i,j,l,m}\phi_{jl}\frac{\partial(f\circ\phi^{-1})}{\partial x_{m}}\phi_{mi}a_{il}\frac{\partial}{\partial x_{j}}$
$=\sum_{i,j}\frac{\partial(f\circ\phi^{-1})}{\partial x_{i}}b_{ij}\frac{\partial}{\partial x_{j}},$
where $(b_{ij})=\phi(a_{ij})\phi^{T}$. Thus, up to a linear change of coordinates, we may assume that the matrix $(a_{ij})$ is in congruence normal form, yielding two possibilities for $v$:
$v_{1}=\frac{\partial f}{\partial x_{1}}\frac{\partial}{\partial x_{0}}-\frac{\partial f}{\partial x_{0}}\frac{\partial}{\partial x_{1}}+\frac{\partial f}{\partial x_{3}}\frac{\partial}{\partial x_{2}}-\frac{\partial f}{\partial x_{2}}\frac{\partial}{\partial x_{3}}\quad\text{or}\quad v_{2}=\frac{\partial f}{\partial x_{1}}\frac{\partial}{\partial x_{0}}-\frac{\partial f}{\partial x_{0}}\frac{\partial}{\partial x_{1}}.$
We will show that none of those cases are possible. By construction,
$0=\iota_{v}(\omega\wedge d\omega)=(\iota_{v}\omega)d\omega-\omega\wedge\iota_{v}d\omega,$
whence $\iota_{v}\omega=0$, otherwise we would get $\omega\wedge d\omega=0$. In particular, we get $\omega\wedge\iota_{v}d\omega=0$, hence $\iota_{v}d\omega=h\omega$ for some degree-one polynomial. In both cases of interest, the vector fields have vanishing divergence. Hence, $d\omega\wedge d\omega=3\iota_{v}(dx_{0}\wedge\dots\wedge dx_{4})$. Moreover,
$0=\iota_{v}(d\omega\wedge d\omega)=2(\iota_{v}d\omega)\wedge d\omega=2h\,\omega\wedge d\omega.$
Therefore, $h=0$ and $\iota_{v}d\omega=0$. We will use this fact to derive a contradiction.
Case 1: Note that the coefficients of $v_{1}$ form a regular sequence of length $4$, hence the associated Koszul complex is exact for polynomial $k$-forms with $2\leq k\leq 4$. Thus $\iota_{v_{1}}d\omega=0$ implies that there exists a homogeneous $3$-form $\theta$ such that $d\omega=\iota_{v}\theta$. However, this is absurd since such $\theta$ would have coefficients of negative degree. Hence, $d\omega=0$ which is also absurd.
Case 2: For $v_{2}$ the equation $\iota_{v_{2}}\omega=0$ implies that
$\omega=\lambda\frac{\partial f}{\partial x_{0}}dx_{0}+\lambda\frac{\partial f}{\partial x_{1}}dx_{1}+a_{2}\,dx_{2}+a_{3}\,dx_{3}+a_{4}\,dx_{4}$
for some $\lambda\in{\mathbb{C}}$ and $a_{j}$ of degree two. Using that $\frac{\partial f}{\partial x_{0}},\frac{\partial f}{\partial x_{1}}$ is a regular sequence, we get from the equation $\iota_{v}d\omega=0$ that, for $i=0,1$ and $j=2,3,4$,
$\frac{\partial a_{j}}{\partial x_{i}}=\lambda\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}.$
Therefore, $\omega=\lambda df+\eta$, where $\eta$ does not depend on $x_{0}$ or $x_{1}$. Note that $0=\iota_{{\rm rad}}\omega=3\lambda\,f+\iota_{{\rm rad}}\eta$, which implies that $\lambda=0$ and $\omega=\eta$. But this is absurd since $\eta$ is integrable. ∎

### v2

@@
 We also prove that the same conclusion holds for any smooth hypersurface of dimension at least three that is not a quadric threefold.
-The proof of these results follows essentially from a more general structure theorem for foliations on manifolds covered by lines.
+The proofs of these results follow essentially from a more general structure theorem for foliations on manifolds covered by lines.
 Key words and phrases:
@@
 1. Introduction
-A singular holomorphic foliation $\mathcal{F}$ of codimension-one on a projective variety $X$ is determined by a $1$-form $\omega$ with values in a line bundle $N$, satisfying the Frobenius integrability condition $\omega\wedge d\omega=0$, and whose singular set has codimension at least two.
+A singular holomorphic foliation $\mathcal{F}$ of codimension one on a projective variety $X$ is determined by a $1$-form $\omega$ with values in a line bundle $N$, satisfying the Frobenius integrability condition $\omega\wedge d\omega=0$, and whose singular set has codimension at least two.
 In this case, $N$ is called the normal bundle of $\mathcal{F}$.
@@
 The irreducible components of these spaces have been studied in various contexts.
+For example, when $X$ is a surface, the integrability condition is automatically satisfied, and the corresponding spaces are irreducible.
 When $\operatorname{Pic}(X)\simeq\mathbb{Z}$, a foliation $\mathcal{F}$ with normal bundle $N$ is said to have degree $d$ if $N\simeq{\mathcal{O}}_{X}(d+2)$, where ${\mathcal{O}}_{X}(1)$ denotes the hyperplane bundle.
 In this case, we denote the space of degree-$d$ foliations by ${\sf Fol}(X,d)$.
-For example, when $X$ is a surface, the integrability condition is automatically satisfied, and the corresponding spaces are irreducible.
-When $X={\mathbb{P}^{n}}$, for $n\geq 3$, the cases $d=0,1,2$ are completely classified, and partial results exist for $d=3$ (see [13], [16], [7], and [10]).
+For $X={\mathbb{P}^{n}}$, $n\geq 3$, the cases $d=0,1,2$ are completely classified, and partial results exist for $d=3$ (see [Deserti-Cerveau], [JP-Jouanoulou79], [LNC-96], and [CLP-2022]).
 For more general smooth projective varieties, few results describe the structure of foliations of low degree.
-When $d=0$, classifications are available for certain homogeneous varieties [5, 19], for hypersurfaces [15], and for smooth complete intersections [1].
+When $d=0$, classifications are available for certain homogeneous varieties [BFM, kuster2025], for hypersurfaces [Fig23], and for smooth complete intersections [ACM-dist].
 In this work, we study the irreducible components of the space of degree-one foliations on smooth projective complete intersections. Our main result is the following:
 Theorem A.
-Let $X$ be a smooth complete intersection in ${\mathbb{P}^{n+s}}$ of type $(d_{1},\ldots,d_{s})$, with $n\geq 4$ and $c_{1}(TX)\geq 2$.
-If either $s=1$ (i.e., $X$ is a hypersurface) or $d_{j}\geq 3$ for all $j=1,\dots,s$, then ${\sf Fol}(X,1)$ has exactly two irreducible components.
+Let $X$ be a smooth complete intersection in $\mathbb{P}^{n+s}$} of type $(d_{1},\ldots,d_{s})$, such that either
+  1. (1)
+    $X$ is a hypersurface, i.e. $s=1$, $n\geq 4$, and $c_{1}(TX)\geq 2$; or
+  2. (2)
+    $d_{j}\geq 3$ for all $j=1,\dots,s$, $n\geq 4$, and $c_{1}(TX)\geq 2$; or
+  3. (3)
+    $d_{j}\geq 4$ for all $j=1,\dots,s$ and $n\geq 3$.
+Then ${\sf Fol}(X,1)$ has exactly two irreducible components.
 The general elements in these components are defined by logarithmic forms (see §2.2).
 In particular, all foliations in these spaces arise as restrictions of foliations on the ambient projective space (see Proposition 3.1).
-When $X$ is the smooth three-dimensional quadric, it is shown in [20] that ${\sf Fol}(X,1)$ possesses an additional irreducible component, consisting of rigid foliations arising from actions of the affine Lie algebra on the quadric, totaling three components.
-Among hypersurfaces, the quadric is the only exceptional case.
-Theorem A applies to all smooth hypersurfaces of degree $\deg(X)\leq n$ and dimension $n\geq 4$.
-Moreover, by [15, Theorem 3.6], every degree-one foliation on a hypersurface $X$ extends to ${\mathbb{P}^{n+1}}$ when $n\geq 3$ and $\deg(X)\geq 4$.
+When $X$ is the smooth three-dimensional quadric, it is shown in [LPT13] that ${\sf Fol}(X,1)$ possesses an additional irreducible component, consisting of rigid foliations arising from actions of the affine Lie algebra on the quadric, totaling three components.
+Among hypersurfaces, the smooth 3-dimensional quadric is the only exceptional case.
+Theorem A applies to all smooth hypersurfaces either of degree $\deg(X)\leq n$ and dimension $n\geq 4$; or degree $\deg(X)\geq 4$ and dimension $n\geq 3$.
 We complement this result by showing that it also holds for cubic threefolds:
@@
 To prove Theorem A, we focus on varieties covered by lines.
-We study the deformations of lines along the leaves of a foliation, as introduced in [21].
+We study the deformations of lines along the leaves of a foliation, as introduced in [LPT-def].
 First, we show that a general free line is not invariant under any foliation (see Proposition 4.7).
@@
 Let $X\subset{\mathbb{P}^{N}}$ be a smooth projective variety of dimension $n\geq 3$ with $\operatorname{Pic}(X)=\mathbb{Z}$, covered by lines.
-If $\mathcal{F}$ is a codimension-one singular holomorphic foliation on $X$ of degree $d$ such that either $d=1$ or
+If $\mathcal{F}$ is a codimension-one holomorphic foliation on $X$ of degree $d$ such that either $d=1$ or
 $c_{1}(TX)\geq n+d-2,$
@@
 In the latter case, we prove that the foliation is the pullback of a degree-one foliation on the projective plane, and is therefore again defined by a closed rational 1-form without codimension-one zeros.
+The last item follows from Theorem 3.2.
 This completes the proof of Theorem A.
-Finally, we apply the same techniques to analyze the structure of degree-two foliations on smooth quadrics (see Corollary 4.10 and Theorem 4.11), providing a partial classification for this setting and laying groundwork for future results.
 This article is organized as follows.
@@
 Let $X$ be a smooth projective complex manifold.
-A distribution $\mathcal{F}$ on $X$ is given by a saturated subsheaf ${T_{\mathcal{F}}}\subset TX$, often called the tangent sheaf of $\mathcal{F}$.
-Here, by saturated we mean that the quotient $TX/{T_{\mathcal{F}}}$ is a torsion-free sheaf.
-The singular locus of $\mathcal{F}$, denoted by $\operatorname{Sing}(\mathcal{F})$, is the locus where the quotient $TX/{T_{\mathcal{F}}}$ is not locally free.
-Since $TX/{T_{\mathcal{F}}}$ is torsion-free, it follows that $\operatorname{codim}\operatorname{Sing}(\mathcal{F})\geq 2$.
-We define the normal sheaf of $\mathcal{F}$ as ${N_{\mathcal{F}}}=(TX/{T_{\mathcal{F}}})^{\vee\vee}$.
+A distribution $\mathcal{F}$ on $X$ is the data of a saturated subsheaf ${T_{\mathcal{F}}}\subset{TX}$, often called the tangent sheaf of $\mathcal{F}$.
+The quotient ${N_{\mathcal{F}}}=TX/{T_{\mathcal{F}}}$ is a torsion-free sheaf called the normal sheaf.
+The singular locus $\operatorname{sing}(\mathcal{F})$ is the locus where $N_{\mathcal{F}}$} is not locally free; since $N_{\mathcal{F}}$} is torsion free, $\operatorname{codim}\operatorname{sing}(\mathcal{F})\geq 2$.
+Over the smooth locus $X\setminus\operatorname{sing}(\mathcal{F})$, these sheaves fit in a short exact sequence:
+$0\longrightarrow{T_{\mathcal{F}}}\longrightarrow{TX}\longrightarrow{N_{\mathcal{F}}}\longrightarrow 0.$
 The tangent sheaf $T_{\mathcal{F}}$} is involutive if it is closed under the Lie bracket of vector fields; in this case, $\mathcal{F}$ defines a foliation, i.e., through every point $x\in X\setminus\operatorname{sing}(\mathcal{F})$ passes an immersed analytic submanifold $\mathcal{L}$ of dimension $\operatorname{rk}{T_{\mathcal{F}}}$ such that $T_{x}\mathcal{L}=T\mathcal{F}_{x}$; these are the leaves of the foliation.
 The map ${TX}\to{N_{\mathcal{F}}}$ is defined by contraction with a twisted $p$-form $\omega_{\mathcal{F}}\in H^{0}(X,\Omega^{p}_{X}\otimes\det({N_{\mathcal{F}}}))$, where $p=\operatorname{rk}{N_{\mathcal{F}}}$.
-Conversely, given a line bundle $L\in\operatorname{Pic}(X)$, $\omega\in H^{0}(X,\Omega^{p}_{X}\otimes L)$ defines a distribution if and only if it is locally decomposable off the singular set, LDS for short, see [11].
+Conversely, given a line bundle $L\in\operatorname{Pic}(X)$, $\omega\in H^{0}(X,\Omega^{p}_{X}\otimes L)$ defines a distribution if and only if it is locally decomposable off the singular set, LDS for short, see [Med].
 This condition means that for every point $x\in X\setminus\operatorname{sing}(\omega)$ there exists an open neighborhood $U\ni x$ and $1$-forms $\alpha_{1},\dots,\alpha_{p}$ such that
@@
 ${\sf Fol}^{p}(X,L)\subset{\mathbb{P}}H^{0}(X,\Omega^{p}_{X}\otimes L).$
-The main objective of our work is to describe the irreducible components of this space for $p=1$, $X\subset{\mathbb{P}^{n+1}}$ a smooth complete intersection and specified $L$.
+The main objective of our work is to describe the irreducible components of this space for $p=1$, $X\subset{\mathbb{P}^{N}}$ a smooth complete intersection, and specified $L$.
 Any line bundle on $\mathbb{P}^{n}$} is isomorphic to ${\mathcal{O}}_{\mathbb{P}^{n}}(k)$ for some $k\in{\mathbb{Z}}$.
@@
 Remark 2.1.
-In [1, 14, 15] the authors study distributions and foliations on smooth complete intersections satisfying an extra hypothesis: $X=V(f_{1},\dots,f_{k})$ with $\deg(f_{1})\leq\dots\leq\deg(f_{k})$ such that $X_{i}=V(f_{1},\dots,f_{i})$ is smooth for every $i$.
+In [ACM-dist, TeseMateus, Fig23] the authors study distributions and foliations on smooth complete intersections satisfying an extra hypothesis: $X=V(f_{1},\dots,f_{k})$ with $\deg(f_{1})\leq\dots\leq\deg(f_{k})$ such that $X_{i}=V(f_{1},\dots,f_{i})$ is smooth for every $i$.
 There are smooth complete intersections without this property; for instance, $X=V(f_{1},f_{2})$ where $\deg(f_{1})<\deg(f_{2})$ and $V(f_{1})$ is singular.
@@
 Note further that a smooth complete intersection only admits distributions and foliations of nonnegative degree.
-Indeed, $H^{0}(X,\Omega_{X}^{p}(d+p+1))=0$ for $d<0$ and $X$ a smooth complete intersection, see [1, Lemma 5.17].
+Indeed, $H^{0}(X,\Omega_{X}^{p}(d+p+1))=0$ for $d<0$ and $X$ a smooth complete intersection, see [ACM-dist, Lemma 5.17].
 2.1. Algebraically integrable foliations
-A foliation on $X$ is algebraically integrable if its general leaf is an algebraic subvariety.
-In this case, there exists a dominant rational map $\pi\colon X\dashrightarrow Y$ onto a projective variety $Y$ such that the closure of a leaf is an irreducible component of a fiber of $\pi$, see [2, §3].
+A foliation on $X$ is algebraically integrable if the leaf through a general point is an algebraic subvariety.
+In this case, there exists a dominant rational map $\pi\colon X\dashrightarrow Y$ onto a projective variety $Y$ such that the closure of a leaf is an irreducible component of a fiber of $\pi$, see [AD-fano, §3].
 Proposition 2.2.
-Let $X$ be a complex projective manifold of dimension $n\geq 3$ with $\operatorname{Pic}(X)\cong{\mathbb{Z}}$.
+Let $X$ be a complex projective manifold of dimension greater than $2$ and with $\operatorname{Pic}(X)\cong{\mathbb{Z}}$.
 Let $\mathcal{F}$ be a codimension-one foliation on $X$.
-Suppose that $\mathcal{F}$ is algebraically integrable, then either:
+Suppose that $\mathcal{F}$ is algebraically integrable.
+Then either:
   1. (1)
@@
 Proof.
-This follows directly from the arguments in [20, Lemma 3.1 and Theorem 3.2]. ∎
+This follows directly from the arguments in [LPT13, Lemma 3.1 and Theorem 3.2]. ∎
 Corollary 2.3.
-Let $X\subset{\mathbb{P}^{N}}$ be a smooth complete intersection of dimension $n\geq 3$ and $\mathcal{F}$ be a codimension-one singular holomorphic foliation on $X$ of degree $d$.
+Let $X\subset{\mathbb{P}^{N}}$ be a smooth complete intersection of dimension $n\geq 3$ and $\mathcal{F}$ be a codimension-one holomorphic foliation on $X$ of degree $d$.
 If $\mathcal{F}$ is algebraically integrable, then
@@
 Proof.
-By Proposition 2.2, either $\mathcal{F}$ admits a rational first integral $\rho\colon X\dashrightarrow{\mathbb{P}^{1}}$ with only irreducible fibers or it admits a subfoliation $\mathcal{G}$.
+By Proposition 2.2, either $\mathcal{F}$ admits a rational first integral $\rho\colon X\dashrightarrow{\mathbb{P}^{1}}$ with only irreducible fibers or admits a subfoliation $\mathcal{G}$.
 Suppose we are in the first case.
 By Lefschetz Hyperplane Theorem, $X$ is simply connected.
-Then the argument follows the proof of [20, Proposition 3.5].
-By [20, Theorem 3.3], $\rho$ admits at most $2$ multiple fibers, which we may assume to be over 00 and $\infty$, and write $\rho^{-1}(0)=pH_{0}$ and $\rho^{-1}(\infty)=qH_{\infty}$.
+Then the argument follows the proof of [LPT13, Proposition 3.5].
+By [LPT13, Theorem 3.3], $\rho$ admits at most $2$ multiple fibers, which we may assume to be over 0 and $\infty$, and write $\rho^{-1}(0)=pH_{0}$ and $\rho^{-1}(\infty)=qH_{\infty}$.
 Then $\rho^{*}(\frac{dx}{x}-\frac{dy}{y})$ is a logarithmic $1$-form defining $\mathcal{F}$ with polar divisor $H_{0}+H_{\infty}$ and empty zero divisor.
 ∎
-Notice that the proof of corollary holds for any simply connected manifold $X$ with Picard number $1$.
+Notice that the proof of Corollary 2.3 holds for any simply connected manifold $X$ with Picard number $1$.
 For our purposes, we will only need $X$ to be a smooth complete intersection, especially hypersurfaces.
@@
 $\omega=L_{1}dL_{2}-L_{2}dL_{1},$
-where $L_{1}$ and $L_{2}$ are degree one homogeneous polynomials, and thus they are all conjugate to each other (see [13, Proposition 3.1]).
+where $L_{1}$ and $L_{2}$ are degree one homogeneous polynomials, and thus they are all conjugate to each other (see [Deserti-Cerveau, Proposition 3.1]).
 Therefore, ${\sf Fol}({\mathbb{P}^{n}},0)$ is isomorphic to the Grassmannian of lines in $\mathbb{P}^{n}$}.
-For codimension-one foliations of degree one, Jouanoulou in [16] proved that ${\sf Fol}({\mathbb{P}^{n}},1)$ has two irreducible components.
+For codimension-one foliations of degree one, Jouanoulou in [JP-Jouanoulou79] proved that ${\sf Fol}({\mathbb{P}^{n}},1)$ has two irreducible components.
 Both components parameterize logarithmic foliations, which we now describe in more generality.
-Let $X$ be any smooth projective variety such that $\dim(X)\geq 3$, $\operatorname{Pic}(X)={\mathbb{Z}}$ and $H^{1}(X,{\mathbb{C}})=0$.
+Let $X$ be any smooth projective complex variety such that $\dim(X)\geq 3$, $\operatorname{Pic}(X)={\mathbb{Z}}$ and $H^{1}(X,{\mathbb{C}})=0$.
 Given $d_{1},d_{2},\dots,d_{k}$ positive integers and $\lambda_{1},\lambda_{2},\dots,\lambda_{k}$ complex numbers, not all zero, such that $\sum\lambda_{i}\cdot d_{i}=0$, we can take sections $f_{i}\in H^{0}(X,{\mathcal{O}}_{X}(d_{i}))$ to form global sections
@@
 Then $\omega$ is integrable and, if its singular locus has codimension greater than one, it determines a foliation known as a logarithmic foliation.
-Varying $\lambda_{i}$ and $f_{i}$, $i=1,\dots,k$, the forms $\omega$ sweep an irreducible algebraic subvariety of ${\sf Fol}(X,d_{1}+\dots+d_{k}-2)$ whose closure is denoted by $\textsf{Log}(X;d_{1},d_{2},\dots,d_{k})$.
-In [6], it was established that $\textsf{Log}(X;d_{1},d_{2},\dots,d_{k})$ is an irreducible component of ${\sf Fol}(X,\sum d_{i}-2)$.
+Varying $\lambda_{i}$ and $f_{i}$, $i=1,\dots,k$, the forms $\omega$ sweep an irreducible algebraic variety subvariety of ${\sf Fol}(X,d_{1}+\dots+d_{k}-2)$ whose closure is denoted by $\textsf{Log}(X;d_{1},d_{2},\dots,d_{k})$.
+In [Calvo94], it was established that $\textsf{Log}(X;d_{1},d_{2},\dots,d_{k})$ is an irreducible component of ${\sf Fol}(X,\sum d_{i}-2)$.
 It is called the logarithmic component.
-If $k=2$, it is also called a rational component.
+When $k=2$, the foliation defined by this 1-form $\omega=\lambda_{1}f_{2}df_{1}+\lambda_{2}f_{1}df_{2}$ is algebraically integrable; this component is also called a rational component.
 These varieties can also be defined by the closure of the rational map
@@
 where $\mathbb{L}_{2}$ is the set of linear projections of $\mathbb{P}^{n}$} to $\mathbb{P}^{2}$}.
-Then, in [8], it was demonstrated that $\textsf{LPB}(d)$ is an irreducible component of ${\sf Fol}({\mathbb{P}^{n}},d)$, and it is called the Linear Pull-Back component.
+Then, in [CL-NE], it was demonstrated that $\textsf{LPB}(d)$ is an irreducible component of ${\sf Fol}({\mathbb{P}^{n}},d)$, and it is called the linear pullback component.
 It is an elementary exercise to show that $\textsf{LPB}(1)=\textsf{Log}({\mathbb{P}^{n}};1,1,1)$.
-For degree two, Cerveau and Lins Neto proved in [7] that ${\sf Fol}({\mathbb{P}^{n}},2)$ has six irreducible components: a pull-back Linear component $\textsf{LPB}(2)$; the rational components $\textsf{Log}({\mathbb{P}^{n}};3,1)$ and $\textsf{Log}({\mathbb{P}^{n}};2,2)$; the logarithmic components $\textsf{Log}({\mathbb{P}^{n}};2,1,1)$ and $\textsf{Log}({\mathbb{P}^{n}};1,1,1,1)$; and another irredutible component called exeptional component, which is determined by the closure of the map
+In degree two, Cerveau and Lins Neto proved in [LNC-96] that ${\sf Fol}({\mathbb{P}^{n}},2)$ has six irreducible components: a linear pullback component $\textsf{LPB}(2)$; the rational components $\textsf{Log}({\mathbb{P}^{n}};3,1)$ and $\textsf{Log}({\mathbb{P}^{n}};2,2)$; the logarithmic components $\textsf{Log}({\mathbb{P}^{n}};2,1,1)$ and $\textsf{Log}({\mathbb{P}^{n}};1,1,1,1)$; and another irredutible component called exeptional component, which is determined by the closure of the map
 $\begin{array}[]{rcl}\Psi\colon\mathbb{L}_{3}&\rightarrow&{\sf Fol}({\mathbb{P}^{n}},2)\\ \pi&\mapsto&\pi^{*}\mathcal{E}\end{array},$
 where $\mathcal{E}$ is the exceptional foliation determined by the section
-$\beta=\frac{1}{x_{0}}(2fdg-3gdf)\in H^{0}({\mathbb{P}^{3}},\Omega^{1}_{{\mathbb{P}^{3}}}(4)).$
-where $f=2x_{0}x_{2}-x_{1}^{2}$ and $g=3x_{0}^{2}x_{3}-3x_{0}x_{1}x_{2}+x_{1}^{3}$.
-In [10], the irreducible components of ${\sf Fol}({\mathbb{P}^{n}},3)$ were partially classified; there are at least 24 irreducible components.
-Besides the projective space, there is some literature on irreducible components of the space of foliations on, for instance, Fano three-folds [20], homogeneous spaces [5, 19], and hypersurfaces [1, 15].
-In the case of smooth projective hypersurfaces $X$ in ${\mathbb{P}^{n+1}}$, $n\geq 3$, if
+$\beta=\frac{1}{x_{0}}(2fdg-3gdf)\in H^{0}({\mathbb{P}^{3}},\Omega^{1}_{{\mathbb{P}^{3}}}(4)),$
+for $f=2x_{0}x_{2}-x_{1}^{2}$ and $g=3x_{0}^{2}x_{3}-3x_{0}x_{1}x_{2}+x_{1}^{3}$.
+In [CLP-2022], the irreducible components of ${\sf Fol}({\mathbb{P}^{n}},3)$ were partially classified; there are at least 24 irreducible components.
+Besides the projective space, there is some literature on irreducible components of the space of foliations on, for instance, Fano three-folds [LPT13], homogeneous spaces [BFM, kuster2025], and hypersurfaces [ACM-dist, Fig23].
+In the case of smooth projective hypersurfaces $X$ in $\mathbb{P}^{n+1}$}, $n>3$, if
 $\deg(X)>2\deg(\mathcal{F})+1,$
-then the restriction of foliations from the ambient ${\mathbb{P}^{n+1}}$ defines an isomorphism between ${\sf Fol}(X,d)$ and ${\sf Fol}({\mathbb{P}^{n+1}},d)$, see [15, Theorem 3.6].
+then the restriction of foliations from the ambient $\mathbb{P}^{n+1}$} defines an isomorphism between ${\sf Fol}(X,d)$ and ${\sf Fol}({\mathbb{P}^{n+1}},d)$, see [Fig23, Theorem 3.6].
 Therefore, the previous classifications for foliations on projective spaces are valid for these varieties.
 The problem is then to study low-degree hypersurfaces; for degree-one foliations, we need to analyze hypersurfaces of degree at most three.
-In [20], the authors proved that for $X\subset{\mathbb{P}^{4}}$ a smooth quadric hypersurface, ${\sf Fol}(X,1)$ has three irreducible components; only two are restrictions from foliations on ${\mathbb{P}^{4}}$.
+In [LPT13], the authors proved that for $X\subset{\mathbb{P}^{4}}$ a smooth quadric hypersurface, ${\sf Fol}(X,1)$ has three irreducible components; only two are restrictions from foliations on $\mathbb{P}^{4}$}.
 In this work, we prove that, for hypersurfaces, this is the only case where ${\sf Fol}(X,1)$ has more than two irreducible components.
 2.3. Degree-zero distributions on the projective space
-In [1], the authors classify distributions of codimension one and degree zero on ${\mathbb{P}^{n}}$.
+In [ACM-dist], the authors classify distributions of codimension one and degree zero on $\mathbb{P}^{n}$}.
 There is a stratification according to the class of the $1$-form defining the distribution.
@@
 Proof.
-Let $\omega\in H^{0}({\mathbb{P}^{n}},\Omega_{\mathbb{P}^{n}}^{q}(q+1))$ a $q$-form defining $\mathcal{D}$.
-By [11, Theorem A], there exists a choice of homogeneous coordinates such that either $\omega=\alpha\wedge dx_{0}\wedge\cdots\wedge dx_{q-2}$ for some $\alpha$ a $1$-form with linear coefficients, or $\omega=i_{v}dx_{0}\wedge\cdots\wedge dx_{q}$, for some linear vector field $v$.
+Let $\omega\in H^{0}({\mathbb{P}^{n}},\Omega_{\mathbb{P}^{n}}^{q}(q+1))$ be a $q$-form defining $\mathcal{D}$.
+By [Med, Theorem A], there exists a choice of homogeneous coordinates such that either $\omega=\alpha\wedge dx_{0}\wedge\cdots\wedge dx_{q-2}$ for some 1-form $\alpha$ with linear coefficients, or $\omega=i_{v}dx_{0}\wedge\cdots\wedge dx_{q}$, for some linear vector field $v$.
 Imposing $i_{\rm rad}\omega=0$, for ${\rm rad}=\sum_{j}x_{j}\frac{\partial}{\partial x_{j}}$ the radial vector field, we must have $\omega=i_{\rm rad}dx_{0}\wedge\cdots\wedge dx_{q}$, up to a scalar multiple.
@@
 3. Extensions and restrictions of foliations on complete intersections
-Let $U_{k}$ be the open subset of twisted $q$-forms in $H^{0}({\mathbb{P}^{n}},\Omega_{\mathbb{P}^{n}}^{q}(k))$ whose singular set has codimension greater than or equal to two and let $X\subset{\mathbb{P}^{n}}$ be a smooth complete intersection.
+Let $U_{k}$ be the open subset of twisted $q$-forms in $H^{0}({\mathbb{P}^{N}},\Omega_{\mathbb{P}^{N}}^{q}(k))$ whose singular set has codimension greater than or equal to two, and let $X\subset{\mathbb{P}^{N}}$ be a smooth complete intersection.
 Definition 3.1.
 We say that $\omega\in U_{k}$ is transversal in codimension one to $X$ if $i^{*}\omega$ has a singular set of codimension greater than or equal to two, where $i\colon X\rightarrow{\mathbb{P}^{N}}$ is the natural inclusion.
-A singular holomorphic foliation $\mathcal{F}$ of codimension $q$ is transversal in codimension one to $X$, if a twisted $q$-form that determines it is transversal in codimension-one to $X$.
-If a foliation $\mathcal{F}$ is transversal in codimension one to $X$, the restriction to $X$ of its twisted $q$-form naturally satisfies the decomposability and integrability hypotheses and, therefore, defines a foliation $\mathcal{G}$ on $X$.
+A singular holomorphic foliation $\mathcal{F}$ on $\mathbb{P}^{N}$} of codimension $q$ is transversal in codimension one to $X$, if a twisted $q$-form that determines it is transversal in codimension one to $X$.
+If a foliation $\mathcal{F}$ on $\mathbb{P}^{N}$} is transversal in codimension one to $X$, the restriction to $X$ of its twisted $q$-form naturally satisfies the decomposability and integrability hypotheses and, therefore, defines a foliation $\mathcal{G}$ on $X$.
 Definition 3.2.
@@
 Therefore, we can find coprime polynomials $F,G$ such that $\operatorname{rest}_{0}(F)=f$ and $\operatorname{rest}_{0}(G)=g$, and the foliation restriction generated by $\deg(g)GdF-\deg(f)FdG$ is the extension of $\mathcal{F}$.
-In general, for logarithmic forms as in (2.1), we have the commutative diagram
-[…]
-which projects all logarithmic forms of ${\mathbb{P}^{n}}$ onto $X$ surjectively, where $i\colon X\rightarrow{\mathbb{P}^{n}}$ is the natural inclusion map.
-Therefore, all foliations generated by logarithmic forms extend.
+In general, for logarithmic 1-forms as in (2.1), we have the commutative diagram
+(3.1) ${\Sigma\times\left(\prod^{k}_{i=1}\mathbb{P}H^{0}({\mathbb{P}^{N}},\mathcal{O}_{{\mathbb{P}^{N}}}(d_{i}))\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\Phi_{{\mathbb{P}^{N}}}}$$\scriptstyle{\text{id}\times\operatorname{rest}_{0}}$${\mathbb{P}H^{0}({\mathbb{P}^{N}},\Omega_{\mathbb{P}^{N}}^{1}(\sum d_{i}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i^{*}}$${\Sigma\times\left(\prod^{k}_{i=1}\mathbb{P}H^{0}(X,{\mathcal{O}}_{X}(d_{i}))\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\Phi_{X}}$${\mathbb{P}H^{0}(X,\Omega_{X}^{1}(\sum d_{i}))}$
+which projects all logarithmic 1-forms of $\mathbb{P}^{N}$} onto $X$ surjectively, where $i\colon X\rightarrow{\mathbb{P}^{N}}$ is the natural inclusion map.
+Therefore, all foliations generated by logarithmic 1-forms extend.
 Proposition 3.1.
@@
 Proof.
-As seen above, if $\mathcal{F}$ is generated by a logarithmic form, it has an extension.
+As seen above, if $\mathcal{F}$ is generated by a logarithmic 1-form, it has an extension.
 Furthermore, the logarithmic and rational components can be determined by the closure of the image of a rational map as (3.1).
 Suppose $\mathcal{F}$ is in a logarithmic component $\textsf{Log}(X,d_{1},d_{2},\dots,d_{k})\subset{\sf Fol}(X,\sum d_{i}-2)$ and is an adherence value of the map $\Phi$.
-In this case, there exists a sequence of foliations $\{\mathcal{F}_{i}\}_{i=0}^{\infty}$ in the logarithmic component such that $\mathcal{F}_{i}\rightarrow\mathcal{F}$ and each $\mathcal{F}_{i}$ is defined by a logarithmic form.
-Since each foliation $\mathcal{F}_{i}$ extends, there exists a sequence of foliations $\{\mathcal{G}_{i}\}_{i=0}^{\infty}$ in the logarithmic component $\textsf{Log}({\mathbb{P}^{n}},d_{1},d_{2},\dots,d_{k})$, where ${\left.\kern-1.2pt\mathcal{G}_{i}\vphantom{\big{|}}\right|_{X}}=\mathcal{F}_{i}$ and each $\mathcal{G}_{i}$ is determined by a logarithmic form $\omega_{i}$.
+In this case, there exists a sequence of foliations $\{\mathcal{F}_{i}\}_{i=0}^{\infty}$ in the logarithmic component such that $\mathcal{F}_{i}\rightarrow\mathcal{F}$ and each $\mathcal{F}_{i}$ is defined by a logarithmic 1-form.
+Since each foliation $\mathcal{F}_{i}$ extends, there exists a sequence of foliations $\{\mathcal{G}_{i}\}_{i=0}^{\infty}$ in the logarithmic component $\textsf{Log}({\mathbb{P}^{N}},d_{1},d_{2},\dots,d_{k})$, where ${\left.\kern-1.2pt\mathcal{G}_{i}\vphantom{\big|}\right|_{X}}=\mathcal{F}_{i}$ and each $\mathcal{G}_{i}$ is determined by a logarithmic 1-form $\omega_{i}$.
 Since the integrability condition determines a proper closed set in the class space $\mathbb{P}H^{0}({\mathbb{P}^{N}},\Omega_{\mathbb{P}^{N}}^{1}(k)),$ the set of points adhering to the sequence $([\omega_{i}])_{i=0}^{\infty}$ is nonempty and each point in this set is integrable.
-Thus, given $[\omega]$ as one of these adherent points, its restriction to $X$ determines $\mathcal{F}$ by definition.
+Thus, given $\omega]$ as one of these adherent points, its restriction to $X$ determines $\mathcal{F}$ by definition.
 Furthermore, we have that $\operatorname{codim}(\omega)\geq 2$, otherwise, $\operatorname{codim}(\operatorname{rest}(\omega))\leq 1$ and $\operatorname{rest}(\omega)$ would not determine $\mathcal{F}$.
@@
 ∎
+In general, foliations on a complete intersection $X$ extend to the projective space when the degree of the foliation is sufficiently small compared to the degrees of the hypersurfaces defining the complete intersection.
+In the case of hypersurfaces, [Fig23, Theorem 3.6] shows that a foliation $\mathcal{F}$ has an unique extension if $\deg(X)>2\deg(\mathcal{F})+1$.
+For degree-one foliations on complete intersections, we have the following result, cf.
+[TeseMateus, Teorema 2.1.1].
+Theorem 3.2.
+Let $X$ be a smooth complete intersection in $\mathbb{P}^{n+s}$} of type $(d_{1},\ldots,d_{s})$ and dimension $n\geq 3$.
+If $d_{j}\geq 4$ for $j=1,\dots,s$, then any codimension-one foliation on $X$ of degree one has a unique extension to $\mathbb{P}^{n+s}$}.
+Proof.
+Let $\omega\in H^{0}(X,\Omega^{1}_{X}(\deg(3))$ be a section determining $\mathcal{F}$ on $X$.
+In light of Remark 2.1, [ACM-dist, Proposition 5.22] holds for smooth complete intersections as in the hypotheses above.
+Thus, the restriction map
+$r_{1}\colon H^{0}({\mathbb{P}^{n+s}},\Omega^{1}_{{\mathbb{P}^{n+s}}}(3))\rightarrow H^{0}(X,\Omega^{1}_{X}(3)$
+is an isomorphism, and
+$r_{3}\colon H^{0}({\mathbb{P}^{n+s}},\Omega^{3}_{{\mathbb{P}^{n+s}}}(6))\rightarrow H^{0}(X,\Omega^{3}_{X}(6))$
+is injective.
+Therefore, there exists a unique section $\alpha\in H^{0}({\mathbb{P}^{n+s}},\Omega^{1}_{{\mathbb{P}^{n+s}}}(3))$ whose restriction to $X$ is $\omega$.
+Note that $\alpha$ must have singularities in codimension at least two, as so does $\omega$.
+Observe that
+$r_{3}(\alpha\wedge d\alpha)=\omega\wedge d\omega=0.$
+Then $\alpha$ is integrable by the injectivity of $r_{3}$. ∎
+An immediate consequence is that the space of degree-one foliations on smooth complete intersections satisfying the hypothesis of the theorem is (set-theoretically) the same as the one for the projective space, cf.
+[TeseMateus, Corolário 2.2.7].
+Corollary 3.3.
+Let $X$ be a smooth complete intersection in $\mathbb{P}^{n+s}$} of type $(d_{1},\ldots,d_{s})$ and dimension $n\geq 3$. If $d_{j}\geq 4$ for $j=1,\dots,s$, then
+${\sf Fol}(X,1)=\textsf{Log}(X;1,2)\cup\textsf{Log}(X;1,1,1).$
 3.2. Codimension two foliations
-For codimension-two foliations of degree zero, we have the following results regarding extensions.
+For codimension-two degree-one foliations, we have the following results regarding extensions.
 Theorem 3.4.
@@
 In particular, $\mathcal{F}$ is a codimension-two distribution on $X$ of degree zero.
-Applying [14, Corolário 2.3.3], we see that $\mathcal{F}$ has a unique extension as a codimension-two distribution to ${\mathbb{P}^{n+s}}$ of degree zero.
+Applying [TeseMateus, Corolário 2.3.3], we see that $\mathcal{F}$ has a unique extension as a codimension-two distribution to $\mathbb{P}^{n+s}$} of degree zero.
 Finally, by Proposition 2.4, we conclude that the extension of $\mathcal{F}$ is integrable.
@@
 Corollary 3.5.
-Let $X$ be a smooth hypersurface of ${\mathbb{P}^{n+1}}$, $n\geq 4$.
-If $\mathcal{F}$ is a codimension-two holomorphic foliation on $X$ of degree zero then $\mathcal{F}$ extends on ${\mathbb{P}^{n+1}}$.
+Let $X$ be a smooth hypersurface of $\mathbb{P}^{n+1}$}, $n\geq 4$. If $\mathcal{F}$ is a codimension-two holomorphic foliation on $X$ of degree zero then $\mathcal{F}$ extends on $\mathbb{P}^{n+1}$}.
 Proof.
-It follows by Theorem 3.2, if $\deg(X)\geq 3$ and, for quadrics, by [3, Proposition 3.18]. ∎
+It follows by Theorem 3.4, if $\deg(X)\geq 3$ and, for quadrics, by [AD-fano2, Proposition 3.18]. ∎
 4. Structure theorem for foliations on complete intersections
@@
 The scheme $\operatorname{Mor}({\mathbb{P}^{1}},X)$ parametrizing morphisms from $\mathbb{P}^{1}$} to $X$ is smooth around a free morphism $f$ and its tangent space is naturally isomorphic to $H^{0}(X,f^{*}{TX})$.
-We refer to [12, 18] for further details.
+We refer to [Deb-book, Kollar-book] for further details.
 A particular result we will need is the following.
 Lemma 4.1.
-Let $f\colon{\mathbb{P}^{1}}\to X$ be a free morphism and $Z\subset X$ be a closed subvariety of codimension at least two.
+Let $X$ be a smooth uniruled projective variety, let $f\colon{\mathbb{P}^{1}}\to X$ be a free morphism, and $Z\subset X$ be a closed subvariety of codimension at least two.
 Then, $f^{\prime}({\mathbb{P}^{1}})\cap Z=\emptyset$ for a general deformation $f^{\prime}$ of $f$.
 Proof.
-This is a particular case of [18, II, Proposition 3.7]. ∎
+This is a particular case of [Kollar-book, II, Proposition 3.7]. ∎
 Lemma 4.2.
@@
 Proof.
-See [12, Proposition 4.14] for proof of the existence of $X^{\rm free}$ over any algebraically closed field of characteristic zero.
+See [Deb-book, Proposition 4.14] for proof of the existence of $X^{\rm free}$ over any algebraically closed field of characteristic zero.
 If the base field is uncountable, then $X^{\rm free}\neq\emptyset$.
@@
 We define the tangential foliation of $\mathcal{F}$ on $M$ by the saturation of the image of this morphism $i$ in $T_{M}$.
-In [21], the authors show how the geometry of $\mathcal{F}_{\rm tang}$ impacts the geometry of $\mathcal{F}$.
+In [LPT-def, kuster2025], the authors show how the geometry of $\mathcal{F}_{\rm tang}$ impacts the geometry of $\mathcal{F}$.
 We recall below two of their results.
 Proposition 4.3.
-[21, Proposition 4.10] Let $X$ be a uniruled smooth projective variety and let $\mathcal{F}$ be a codimension one foliation on $X$.
+[LPT-def, Proposition 4.10] Let $X$ be a smooth uniruled projective variety and let $\mathcal{F}$ be a codimension-one foliation on $X$.
 If all leaves of $\mathcal{F}_{\rm tang}$ are algebraic, then there exists an algebraically integrable foliation $\mathcal{G}$ contained in $\mathcal{F}$ such that $H^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})=H^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{G}}})$ for any sufficiently general $f\in M$.
@@
 Theorem 4.4.
-[21, Theorem C] Suppose that $X$ is simply connected, $\operatorname{codim}\mathcal{F}=1$, the general leaf of $\mathcal{F}_{\rm tang}$ is not algebraic, and a general $f\in M$ nontrivially and transversely all the algebraic hypersurfaces invariant by $\mathcal{F}$.
+[LPT-def, Theorem C] Let $X$ be a simply connected uniruled projective manifold and let $\mathcal{F}$ be a codimension-one foliation on $X$.
+Suppose that the general leaf of $\mathcal{F}_{\rm tang}$ is not algebraic, and a general $f]\in M$ intersects nontrivially and transversely all the algebraic hypersurfaces invariant by $\mathcal{F}$.
 Then $\mathcal{F}$ is defined by a closed rational $1$-form without divisorial components in its zero set.
@@
 Since $f^{\prime}$ is tangent to $\mathcal{F}$, we get that $f^{\prime*}\nabla^{B}$ is a holomorphic connection on the vector bundle $f^{\prime*}{N_{\mathcal{F}}}$ on $\mathbb{P}^{1}$}.
-By Weil’s Theorem [4, Theorem 10], $f^{\prime*}{N_{\mathcal{F}}}$ must be trivial.
+By Weil’s Theorem [Atiyah-conn, Theorem 10], $f^{\prime*}{N_{\mathcal{F}}}$ must be trivial.
 Using that $\operatorname{Pic}(X)\cong{\mathbb{Z}}$, we conclude that $c_{1}({N_{\mathcal{F}}})=0$, hence $\mathcal{F}$ is given by a global holomorphic $p$-form $\omega\in H^{0}(X,\Omega^{p}_{X})$, $p=\operatorname{codim}\mathcal{F}$.
-On the other hand, $X$ is uniruled with $\operatorname{Pic}(X)\cong{\mathbb{Z}}$ implies that $X$ is Fano, hence $h^{0}(X,\Omega^{p}_{X})=h^{0,p}(X)=0$, which is a contradiction.
+On the other hand, since $X$ is uniruled with $\operatorname{Pic}(X)\cong{\mathbb{Z}}$, it follows that $X$ is Fano.
+Hence $h^{0}(X,\Omega^{p}_{X})=h^{0,p}(X)=0$, which is a contradiction.
 ∎
@@
 Suppose $X\subset{\mathbb{P}^{N}}$ is a complete intersection given by equations of degrees $d_{1},\dots,d_{s}$.
-If $d_{1}+\cdots+d_{s}\leq N-1$ then $X$ is covered by lines, see [12, Proposition 2.13].
+If $d_{1}+\cdots+d_{s}\leq N-1$ then $X$ is covered by lines, see [Deb-book, Proposition 2.13].
 Then, the previous corollary applies.
@@
 Then, we will apply it to smooth complete intersections and, in particular, to smooth hypersurfaces.
-The result below is inspired by the structure theorem for degree-three foliations on ${\mathbb{P}^{3}}$ by da Costa, Lizarbe, and Pereira, see [10, Theorem A].
+The result below is inspired by the structure theorem for degree-three foliations on $\mathbb{P}^{3}$} by da Costa, Lizarbe, and Pereira, see [CLP-2022, Theorem A].
 Theorem 4.8.
@@
 (4.1) $c_{1}(TX)\geq n+d-2,$
-then one of the following holds:
+then, one of the following holds:
   1. (1)
-    $\mathcal{F}$ is defined by a closed rational 1-form without Codimension one zeros; or
+    $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros; or
   2. (2)
@@
     $\mathcal{F}$ admits an algebraically integrable subfoliation of codimension 2, and degree $k$ such that $k\leq d-2$.
-Note that the inequality $c_{1}(TX)\geq n+d-2$ for $d\geq 2$ only holds for the quadric hypersurface or ${\mathbb{P}^{n}}$, due to the Kobayashi-Ochiai Theorem [17].
+Note that the inequality $c_{1}(TX)\geq n+d-2$ for $d\geq 2$ only holds for the quadric hypersurface or $\mathbb{P}^{n}$}, due to the Kobayashi-Ochiai Theorem [KobOch, Theorem 1.1 and Theorem 2.1].
 Proof.
@@
 $f^{*}TX={\mathcal{O}}_{{\mathbb{P}^{1}}}(2)\oplus{\mathcal{O}}_{{\mathbb{P}^{1}}}(1)^{a}\oplus{\mathcal{O}}_{{\mathbb{P}^{1}}}^{b}$
-where $a=c_{1}(TX)-2\geq 0$ and $b=n-c_{1}(TX)+1\geq 0$ see [12, Proposition 2.13 (b)].
+where $a=c_{1}(TX)-2\geq 0$ and $b=n-c_{1}(TX)+1\geq 0$ see [Deb-book, Proposition 2.13 (b)].
 Since $f$ is generically transverse to $\mathcal{F}$, every summand of the decomposition of $f^{*}{T_{\mathcal{F}}}$ satisfies $a_{i}\leq 1$.
@@
     $f^{*}{T_{\mathcal{F}}}\simeq{\mathcal{O}}_{\mathbb{P}^{1}}(1)^{r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}^{\oplus n-2-r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}(-c)$, for some nonnegative integer $r$ and $c=1$ or $c=2$.
-Suppose that $\mathcal{F}$ is not algebraically integrable.
-We now divide the proof into the cases whether $\mathcal{F}_{\rm tang}$ is algebraically integrable or not.
-If $\mathcal{F}_{\rm tang}$ is not algebraically integrable, then by [21, Theorem C], $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros.
-If $\mathcal{F}_{\rm tang}$ is algebraically integrable, then, as argued in the proof of [21, Theorem A], there exists an algebraically integrable subfoliation $\mathcal{G}\subset\mathcal{F}$ of degree $k$ and codimension $q\geq 2$ such that $h^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{G}}})=h^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})$.
+If $\mathcal{F}_{\rm tang}$ is not algebraically integrable, then by [LPT-def, Theorem C], $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros.
+If $\mathcal{F}_{\rm tang}$ is algebraically integrable and $\mathcal{F}$ is not algebraically integrable, then, as argued in the proof of [LPT-def, Theorem A], there exists an algebraically integrable subfoliation $\mathcal{G}\subset\mathcal{F}$ of degree $k$ and codimension $q\geq 2$ such that $h^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{G}}})=h^{0}({\mathbb{P}^{1}},f^{*}{T_{\mathcal{F}}})$.
 It follows that $f^{*}{T_{\mathcal{F}}}\simeq{\mathcal{O}}_{\mathbb{P}^{1}}(1)^{r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}^{\oplus n-2-r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}(-c)$ and $f^{*}{T_{\mathcal{G}}}={\mathcal{O}}_{\mathbb{P}^{1}}(1)^{r}\oplus{\mathcal{O}}_{\mathbb{P}^{1}}^{\oplus n-2-r}$.
@@
 $c_{1}({T_{\mathcal{F}}})=r-c=c_{1}(TX)-(d+2)$, and $c_{1}({T_{\mathcal{G}}})=r=c_{1}(TX)-(k+3)$.
-Combining and simplifying, we get $k=d-1-c$. This proves the theorem. ∎
+Combining and simplifying, we get k=d1ck=d-1-c. This proves the theorem. ∎
 The previous theorem allows us to determine the structure of degree-one foliations on smooth complete intersections.
 Corollary 4.9.
-Let $X$ be a smooth complete intersection of ${\mathbb{P}^{N}}$ of dimension $n\geq 3$. Suppose $c_{1}(TX)\geq 2$. If $\mathcal{F}$ is a codimension-one foliation on $X$ of degree one, then
+Let $X$ be a smooth complete intersection on $\mathbb{P}^{N}$} of dimension $n\geq 3$. Suppose $c_{1}(TX)\geq 2$. If $\mathcal{F}$ is a codimension-one foliation on $X$ of degree one, then
   1. (1)
@@
 Proof.
-Note that $X$ is covered by lines, see Proposition 4.7.
-By Theorem 4.8, either $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros, is algebraically integrable or has a codimension-two subfoliation of negative degree.
+Let $X\subset{\mathbb{P}^{N}}$ be a smooth complete intersection defined by equations of degrees $d_{1},\dots,d_{s}$.
+Since, by hypothesis, $c_{1}(TX)\geq 2$, we have $d_{1}+\cdots+d_{s}\leq N-1$ and then $X$ is covered by lines, see [Deb-book, Proposition 2.13].
+By Theorem 4.8, either $\mathcal{F}$ is defined by a closed rational 1-form without codimension-one zeros, is algebraically integrable, or admits a codimension-two subfoliation of negative degree.
 The latter is not possible, as observed in Remark 2.2.
-Suppose $\mathcal{F}$ is not given by a closed rational 1-form without codimension-one zeros.
-Then, by Corollary 2.3, $\mathcal{F}$ must have a codimension-two subfoliation of degree zero.
+Assume now that $\mathcal{F}$ is not given by a closed rational 1-form without codimension-one zeros.
+Then, by Corollary 2.3, $\mathcal{F}$ must admit a codimension-two subfoliation of degree zero.
 ∎
 Corollary 4.10.
-Let $\mathcal{F}$ be a codimension-one singular holomorphic foliation of degree two on a smooth quadric ${Q^{n}}\subset{\mathbb{P}^{n+1}}$, $n\geq 4$. Then
+Let $\mathcal{F}$ be a codimension-one holomorphic foliation of degree two on a smooth quadric ${Q^{n}}\subset{\mathbb{P}^{n+1}}$, $n\geq 4$. Then
   1. (1)
@@
 Combining Theorem 4.8 and Corollary 2.3, we have that $\mathcal{F}$ admits a subfoliation $\mathcal{G}$ of codimension two and degree $k\leq 1$.
-In case $k=0$, by [3, Proposition 3.18], $\mathcal{G}$ must be given by a linear projection to ${\mathbb{P}^{2}}$, hence $\mathcal{F}$ is a linear pullback of a foliation of degree two on ${\mathbb{P}^{2}}$.
+In case $k=0$, by [AD-fano2, Proposition 3.18], $\mathcal{G}$ must be given by a linear projection to $\mathbb{P}^{2}$}, hence $\mathcal{F}$ is a linear pullback of a foliation of degree two on $\mathbb{P}^{2}$}.
 ∎
 Theorem 4.11.
-Let $\mathcal{F}$ be a codimension-one singular holomorphic foliation of degree two on a smooth tridimensional quadric ${Q^{3}}$. Then
+Let $\mathcal{F}$ be a codimension-one holomorphic foliation of degree two on a smooth tridimensional quadric $Q^{3}$}. Then
   1. (1)
@@
 The proof is similar to the proof of Corollary 4.10.
-Except for the subfoliation $\mathcal{G}$ has dimension one, and [3, Proposition 3.18] does not apply.
-Instead, ${T_{\mathcal{G}}}={\mathcal{O}}_{Q^{3}}$ and $\mathcal{G}$ is defined by a global vector field on ${Q^{3}}$ without zeros in Codimension-one.
+Except when the subfoliation $\mathcal{G}$ has dimension one, and [AD-fano2, Proposition 3.18] does not apply.
+Instead, ${T_{\mathcal{G}}}={\mathcal{O}}_{Q^{3}}$ and $\mathcal{G}$ is defined by a global vector field on $Q^{3}$} without zeros in codimension one.
 Let $\omega\in H^{0}({Q^{3}},\Omega_{{Q^{3}}}^{1}(3))$ be a 1-form defining $\mathcal{F}$ and define
@@
 Note that $\mathcal{G}$ is defined by a vector field $v\in\mathfrak{fix}(\mathcal{F})$.
-By [20, Lemma 4.3], either $\mathfrak{fix}(\mathcal{F})\neq\mathfrak{aut}(\mathcal{F})$ and $\mathcal{F}$ is defined by a closed rational 1-form without zeros in Codimension-one; or $\mathfrak{fix}(\mathcal{F})=\mathfrak{aut}(\mathcal{F})$ and there exists a non-trivial algebraic action of a one-dimensional Lie group tangent to $\mathcal{F}$; on ${Q^{3}}$ this group must be $\mathbb{C}$ or $\mathbb{C}^{*}$.
-Suppose we are in the latter case and assume $\mathcal{G}$ is defined by said algebraic action.
-Finally, $\mathcal{G}$ has only zero-dimensional singularities, then [20, Lemma 4.2] implies that ${T_{\mathcal{F}}}={\mathcal{O}}_{Q^{3}}\oplus{\mathcal{O}}_{Q^{3}}(-1)$. ∎
+By [LPT13, Lemma 4.3], either $\mathfrak{fix}(\mathcal{F})\neq\mathfrak{aut}(\mathcal{F})$ and $\mathcal{F}$ is defined by a closed rational 1-form without zeros in codimension one; or $\mathfrak{fix}(\mathcal{F})=\mathfrak{aut}(\mathcal{F})$ and there exists a non-trivial algebraic action of a one-dimensional Lie group tangent to $\mathcal{F}$; on $Q^{3}$} this group must be $\mathbb{C}$ or $\mathbb{C}^{*}$.
+Suppose we are in the latter case, and let $\mathcal{G}$ denote the foliation defined by this algebraic action.
+Finally, if $\mathcal{G}$ has only zero-dimensional singularities, then [LPT13, Lemma 4.2] implies that ${T_{\mathcal{F}}}={\mathcal{O}}_{Q^{3}}\oplus{\mathcal{O}}_{Q^{3}}(-1)$.
+∎
 5. Irreducible components of degree-one foliations
-In this section, we apply the structure Theorem 4.8 and its corollaries to determine the irreducible components of ${\sf Fol}(X,1)$ for $X$ either a hypersurface or smooth complete intersection of hypersurfaces of degrees at least three, improving [14, Corolário 2.2.7].
+In this section, we apply the structure Theorem 4.8 and its corollaries to determine the irreducible components of ${\sf Fol}(X,1)$ for $X$ either a hypersurface or smooth complete intersection of hypersurfaces of degrees at least three, improving Corollary 3.3.
 We start with a technical result.
@@
 First, note that $\pi|_{X}$ is surjective.
-In fact, as $12\leq q\leq n-$ 1, we have that the fibers of $\pi$ have dimension at least two, and the intersection of each of these fibers with $X$ is certainly not empty.
+In fact, as $2\leq q\leq n-1$, we have that the fibers of $\pi$ have dimension at least two, and the intersection of each of these fibers with $X$ is certainly not empty.
 Let $\Theta\in H^{0}({\mathbb{P}^{q}},\Omega^{q}_{{\mathbb{P}^{q}}}(q+1))$ be a $q$-twisted form in $\mathbb{P}^{q}$} without zeros; this is possible since $\Omega^{q}_{{\mathbb{P}^{q}}}(q+1)\cong{\mathcal{O}}_{\mathbb{P}^{q}}$.
@@
 Let $\mathcal{F}$ be a codimension-one foliation of degree one on $X$.
-By Corollary 4.9, we have that $\mathcal{F}$ is either defined by a closed rational 1-form without codimension one zeros or $\mathcal{F}$ admits a subfoliation of codimension 2, and degree zero.
-In the first case, $\mathcal{F}$ belongs to $\textsf{Log}(X;1,2)$ or $\textsf{Log}(X;1,1,1)$, see [20, Lemma 2.5].
+By Corollary 4.9, we have that $\mathcal{F}$ is either defined by a closed rational 1-form without codimension-one zeros or $\mathcal{F}$ admits a subfoliation of codimension 2, and degree zero.
+In the first case, $\mathcal{F}$ belongs to $\textsf{Log}(X;1,2)$ or $\textsf{Log}(X;1,1,1)$, see [LPT13, Lemma 2.5].
 Suppose there exists a subfoliation $\mathcal{G}\subset\mathcal{F}$ of codimension two and degree zero.
 By Theorem 3.4, and Corollary 3.5 in the case of hypersurfaces, we have that $\mathcal{G}$ extends to $\mathbb{P}^{n+s}$} and thus $\mathcal{G}$ is defined by the restriction to $X$ of a linear projection $\pi\colon{\mathbb{P}^{n+s}}\dashrightarrow{\mathbb{P}^{2}}$.
-By [9, Lemma 3.1], there exists a foliation $\mathcal{H}$ on ${\mathbb{P}^{2}}$ such that $\mathcal{F}=({\left.\kern-1.2pt\pi\vphantom{\big{|}}\right|_{X}})^{*}\mathcal{H}$.
+By [CLLPT7, Lemma 3.1], there exists a foliation $\mathcal{H}$ on $\mathbb{P}^{2}$} such that $\mathcal{F}=({\left.\kern-1.2pt\pi\vphantom{\big|}\right|_{X}})^{*}\mathcal{H}$.
 Furthermore, Lemma 5.1 guarantees that ${N_{\mathcal{F}}}=({\left.\kern-1.2pt\pi\vphantom{\big|}\right|_{X}})^{*}{N_{\mathcal{H}}}$.
@@
 ∎
+Therefore, Theorem A follows from Theorem 5.2 and Corollary 3.3.
 Note that for $X\subset{\mathbb{P}^{n+1}}$ a hypersurface, $c_{1}(TX)=n+2-\deg(X)$.
 Thus, Theorem 5.2 applies when $\deg(X)\leq n$.
-On the other hand, [15, Theorem 3.6] shows that every degree-one foliation on $X$ extends to ${\mathbb{P}^{n+1}}$ if $\deg(X)\geq 4$ and $n\geq 3$.
-The following is the immediate.
+On the other hand, Corollary 3.3 applies to $\deg(X)\geq 4$ and $n\geq 3$.
+The following is immediate.
 Corollary 5.3.
@@
 This follows by Corollary 5.3 and Proposition 3.1. ∎
-To complete the classification of ${\sf Fol}(X,1)$ for any smooth projective hypersurface $X$ of dimension $n\geq 3$, we only need to analyze the case of cubic hypersurfaces of dimension three.
-The case of a quadric threefold was described in [20, Theorem 5.2].
+To complete the proof of Theorem B about the classification of ${\sf Fol}(X,1)$ for any smooth projective hypersurface $X$ of dimension $n\geq 3$, we only need to analyze the case of cubic hypersurfaces of dimension three.
+The case of a quadric threefold was described in [LPT13, Theorem 5.2].
 In addition to the logarithmic components, there is an irreducible component whose general element is given by an action of the affine group $\mathfrak{aff}(\mathbb{C})$.
@@
 Let $X\subset{\mathbb{P}^{4}}$ be a smooth cubic hypersurface and $\mathcal{F}$ a codimension-one degree-one foliation on $X$ given by $\omega_{0}\in H^{0}(X,\Omega_{X}^{1}(3))$.
-This form extends to (a unique!) $\omega\in H^{0}({\mathbb{P}^{4}},\Omega_{\mathbb{P}^{4}}^{1}(3))$ defining a degree-one distribution on ${\mathbb{P}^{4}}$.
+This form extends to (a unique!) 1-form $\omega\in H^{0}({\mathbb{P}^{4}},\Omega_{\mathbb{P}^{4}}^{1}(3))$ defining a degree-one distribution on $\mathbb{P}^{4}$}.
 If $\omega$ is integrable then $\mathcal{F}$ is the restriction of a degree-one foliation on $\mathbb{P}^{4}$}, thus $\mathcal{F}$ belongs to either $\textsf{Log}(X;2,1)$ or $\textsf{Log}(X;1,1,1)$.